### v1

# Vizing’s Theorem in Near-Linear Time

Abstract
Vizing’s theorem states that any $n$-vertex $m$-edge graph of maximum degree $\Delta$ can be edge colored using at most $\Delta+1$ different colors [Vizing, 1964]. Vizing’s original proof is algorithmic and shows that such an edge coloring can be found in $O(mn)$ time. This was subsequently improved to $\tilde{O}(m\sqrt{n})$ time, independently by [Arjomandi, 1982] and by [Gabow et al., 1985]. [^1] Throughout, we use $\tilde{O}(f):=O(f\operatorname*{polylog}{(n)})$ to suppress log-factors in the number of vertices of the graph.
Very recently, independently and concurrently, using randomization, this runtime bound was further improved to $\tilde{O}(n^{2})$ by [Assadi, 2024] and $\tilde{O}(mn^{1/3})$ by [Bhattacharya, Carmon, Costa, Solomon and Zhang, 2024] (and subsequently to $\tilde{O}(mn^{1/4})$ time by [Bhattacharya, Costa, Solomon and Zhang, 2024]).
We present an algorithm that computes a $(\Delta+1)$-edge coloring in $\tilde{O}(m)$ time—in fact, even $O(m\log{\Delta})$ time—with high probability, giving a near-optimal algorithm for this fundamental problem.
1 Introduction
Given a simple undirected graph $G=(V,E)$ on $n$ vertices and $m$ edges, as well as an integer $\kappa\in\mathbb{N}^{+}$, a $\kappa$-edge coloring $\chi:E\rightarrow\{1,2,\ldots,\kappa\}$ of $G$ assigns a color $\chi(e)$ to each edge $e\in E$ so that any two adjacent edges receive distinct colors. The minimum possible value of $\kappa$ for which a $\kappa$-edge coloring exists in $G$ is known as the edge chromatic number of $G$. If $G$ has maximum vertex degree $\Delta$, any proper edge coloring would require at least $\Delta$ different colors. A classical theorem by Vizing shows that $\Delta+1$ colors are always sufficient [Viz64]. Moreover, it was proven by [Hol81] that it is NP-complete to distinguish whether the edge chromatic number of a given graph is $\Delta$ or $\Delta+1$, and therefore $\Delta+1$ is the best bound we can hope for with polynomial time algorithms.
Vizing’s original proof easily extends to an $O(mn)$ time algorithm, which was improved to $\tilde{O}(m\sqrt{n})$ in the 1980s by Arjomandi [Arj82] and Gabow et al. [GNK+85] independently. More recently, Sinnamon [Sin19] simplified the algorithms of [Arj82, GNK+85] by using randomization, while shaving off extra logarithmic factors from their runtime complexities, achieving a clean $O(m\sqrt{n})$ runtime bound. Very recently, this longstanding $O(m\sqrt{n})$ time barrier was bypassed in two recent concurrent works by Assadi [Ass24] and Bhattacharya et al. [BCC+24] which improve the runtime bound to two incomparable bounds of $\tilde{O}(n^{2})$ and $\tilde{O}(mn^{1/3})$, respectively. In a follow-up work, Bhattacharya et al. [BCSZ25] improved the $\tilde{O}(mn^{1/3})$ time bound of [BCC+24] to $\tilde{O}(mn^{1/4})$.
In this work, we resolve the time complexity of randomized $(\Delta+1)$-edge coloring (up to small logarithmic factors) by presenting a near-linear time algorithm for this problem.
Theorem 1.1.
There is a randomized algorithm that, given any simple undirected graph $G=(V,E)$ on $n$ vertices and $m$ edges with maximum degree $\Delta$, finds a $(\Delta+1)$-edge coloring of $G$ in $\tilde{O}(m)$ time with high probability.
Remarks:
  1. 1.
    Our algorithm can be implemented to run in $O(m\log{\Delta})$ time with high probability. Nonetheless, our main focus is on polynomial runtime improvements in general, and achieving a near-linear time algorithm in particular. For this reason as well as for the sake of simplicity and transparency of our techniques, the main body of the paper presents and analyzes an algorithm that runs in $O(m\log^{3}n)$ time with high probability, which showcases our most novel ideas. A more careful implementation of the same exact algorithm, with proper adaptations, achieves $O(m\log\Delta)$ time; we defer the details of this part to Appendix A.
  2. 2.
    The $O(m\log\Delta)$ time bound that our final algorithm achieves matches the longstanding time bound for $\Delta$-edge coloring bipartite graphs [COS01] (which, to the best of our knowledge, is also the best known runtime for $(\Delta+1)$-edge coloring bipartite graphs). Our algorithm is also related to a recent line of work in [BD23, BD24, Dha24] that focuses on the $\Delta=n^{o(1)}$ case, which in particular gives a randomized $(\Delta+1)$-coloring that runs in $O(m\Delta^{4}\log\Delta)$ time with high probability.
1.1 Related Work
There has been a long line of work on fast algorithms for edge coloring that use more than $\Delta+1$ colors. It was first shown in [KS87] that an edge coloring can be computed in $\tilde{O}(m)$ time when we have $\Delta+\tilde{O}(\sqrt{\Delta})$ different colors. In addition, there are algorithms which run in linear or near-linear time for $(1+\epsilon)\Delta$-edge coloring [DHZ19, BCPS24c, EK24, BD24, Dha24] when $\epsilon\in(0,1)$ is a constant. Most recently, it was shown in [Ass24] that even a $(\Delta+O(\log n))$-edge coloring can be computed in near-linear time.
There are other studies on restricted graph classes. In bipartite graphs, a $\Delta$-edge coloring can be computed in $\tilde{O}(m)$ time [CH82, COS01]. In bounded degree graphs, one can compute a $(\Delta+1)$-edge coloring in $\tilde{O}(m\Delta)$ time [GNK+85], and it was generalized recently for bounded arboricity graphs [BCPS24b]; see also [BCPS24a, CRV24, Kow24] for further recent results on edge coloring in bounded arboricity graphs. Subfamilies of bounded arboricity graphs, including planar graphs, bounded tree-width graphs and bounded genus graphs, were studied in [CY89, CN90, CK08].
Beside the literature in the classical sequential model, considerable effort has been devoted to the study of edge coloring in various computational models in the past few years, including dynamic [BM17, BCHN18, DHZ19, Chr23, BCPS24c, Chr24], online [CPW19, BGW21, SW21, KLS+22, BSVW24, DGS24], distributed [PR01, EPS14, FGK17, GKMU18, BBKO22, CHL+20, Ber22, Chr23, Dav23], and streaming [BDH+19, SB24, CMZ24, GS24] models, among others.
1.2 Technical Overview
We now present an overview of prior approaches to $(\Delta+1)$-edge coloring and describe our techniques at a high level. For the following discussions, we will assume basic familiarity with Vizing’s proof and its underlying algorithm; see Section 2.2 for more details on Vizing’s algorithm.
Prior Approaches: A generic approach for $(\Delta+1)$-edge coloring, dating back to Vizing’s proof itself, is to extend a partial $(\Delta+1)$-edge coloring of the graph one edge at a time, possibly by recoloring some subset of edges, until the entire graph becomes colored. As expected, the main bottleneck in the runtime of this approach comes from extending the coloring to the last few uncolored edges. For instance, given a graph $G=(V,E)$ with maximum degree $\Delta$, we can apply Eulerian partition to divide $G$ into two edge-disjoint subgraphs $G=G_{1}\cup G_{2}$ with maximum degrees at most $\left\lceil\Delta/2\right\rceil$. We then find $(\left\lceil\Delta/2\right\rceil+1)$-edge colorings of the subgraphs $G_{1}$ and $G_{2}$ recursively. Directly combining the colorings of $G_{1}$ and $G_{2}$ gives a coloring of $G$ with $\Delta+3$ colors, so we have to uncolor a $2/(\Delta+3)$ fraction of the edges—amounting to $O(m/\Delta)$ edges—and try to extend the current partial $(\Delta+1)$-edge coloring to these edges. Thus, the “only” remaining part is to figure out a way to color these final $O(m/\Delta)$ edges, and let the above approach take care of the rest.
This task of extending the coloring to the last $\Theta(m/\Delta)$ edges is the common runtime bottleneck of all previous algorithms. Vizing’s original algorithm [Viz64] gives a procedure to extend any partial coloring to an arbitrary uncolored edge $(u,v)$ by rotating some colors around $u$ and flipping the colors of an alternating path starting at $u$. The runtime of this procedure would be proportional to the size of the rotation, usually called a Vizing fan, and the length of the alternating path, usually called a Vizing chain, which are bounded by $\Delta$ and $n$ respectively. As such, the total runtime for coloring the remaining $O(m/\Delta)$ edges using Vizing fans and Vizing chains will be $O(mn/\Delta)$ time.
As one can see, flipping long alternating paths is the major challenge in Vizing’s approach. To improve this part of the runtime, [GNK+85] designed an algorithm that groups all uncolored edges into $O(\Delta^{2})$ types depending on the two colors of the Vizing chain induced by this edge. Since all the Vizing chains of the same type are vertex-disjoint, they can be flipped simultaneously and their total length is only $O(n)$. This means that the runtime of coloring all edges of a single type can be bounded by $O(n)$ time as well. This leads to an $O(n\Delta^{2})$ time algorithm for handling all $O(\Delta^{2})$ types; a more careful analysis can bound this even with $O(m\Delta)$ time. Finally, balancing the two different bounds of $O(mn/\Delta)$ and $O(m\Delta)$ yields a runtime bound of $O(m\sqrt{n})$ for coloring $O(m/\Delta)$ edges, which leads to an $\tilde{O}(m\sqrt{n})$ time algorithm using the above framework.
There has been some very recent progress that broke through this classical barrier of $O(m\sqrt{n})$ time in [GNK+85, Arj82]. In [BCC+24], the authors proved a new useful condition for extending the coloring to uncolored edges: if the uncolored edges admit a small vertex cover, then there is also a faster algorithm for extending the coloring to them. They then show how to precondition the problem so that uncolored edges do indeed admit a small vertex cover, leading to an algorithm with $\tilde{O}(mn^{1/3})$ time. In [Ass24], the author avoided the need for Eulerian partition and recursion altogether by instead designing a new near-linear time algorithm for $(\Delta+O(\log n))$-edge coloring. This algorithm borrows insights from sublinear matching algorithms in regular bipartite graphs by [GKK10] and is thus completely different from other edge coloring algorithms mentioned above. By using this algorithm, finding a $(\Delta+1)$-edge coloring directly reduces to the color extension problem with $O((m\log n)/\Delta)$ uncolored edges (by removing the colors of a $\Theta(\log{n})/\Delta$ fraction of the edges in the $(\Delta+O(\log{n}))$-edge coloring to obtain a partial $(\Delta+1)$-edge coloring first). Applying Vizing’s procedure for these uncolored edges takes additional $\tilde{O}(mn/\Delta)=\tilde{O}(n^{2})$ time, leading to an $\tilde{O}(n^{2})$ time algorithm for $(\Delta+1)$-edge coloring. Finally, in [BCSZ25], the authors showed that a $(\Delta+1)$-coloring can be computed in $\tilde{O}(mn^{1/4})$ time, by using the algorithm of [Ass24] for initial coloring of the graph and then presenting an improved color extension subroutine with a runtime of $\tilde{O}(\Delta^{2}+\sqrt{\Delta n})$ for coloring each remaining edge; the best previous color extension time bounds were either the trivial $O(n)$ bound or the bound $\tilde{O}(\Delta^{6})$ by [Ber22].
Our Approach: A Near-Linear Time Color Extension Algorithm
Our algorithm takes an entirely different path compared to previous work on this problem; specifically, we no longer attempt to design a faster color extension for a single edge, and instead color them in large batches, which allows for a much better amortization of runtime in coloring multiple edges. This ultimately leads to our main contribution — a new randomized algorithm for solving the aforementioned color extension problem for the last $O(m/\Delta)$ edges in $\tilde{O}(m)$ time. With this algorithm at hand, we can simply follow the Eulerian partition approach mentioned earlier and obtain an algorithm for $(\Delta+1)$-edge coloring whose runtime $T(m)$ follows the recursion $T(m)\leq 2T(m/2)+\tilde{O}(m)$; this implies that $T(m)=\tilde{O}(m)$, hence, giving us a near-linear time algorithm for $(\Delta+1)$-edge coloring. This way, we will not even need to rely on the $(\Delta+O(\log{n}))$-edge coloring algorithm of [Ass24] to color the earlier parts of the graph (although one can use that algorithm instead of Eulerian partition approach to the same effect).
We now discuss the main ideas behind our color extension algorithm. In the following, it helps to think of the input graph as being near-regular (meaning that the degree of each vertex is $\Theta(\Delta)$), and thus the total number of edges will be $m=\Theta(n\Delta)$; this assumption is not needed for our algorithm and is only made here to simplify the exposition.
Color Type Reduction: Recall that the runtime of $O(n\Delta^{2})$ for the color extension algorithm of [GNK+85] is due to the fact that there are generally $O(\Delta^{2})$ types of alternating paths in the graph and that the total length of the paths of each color type is bounded by $O(n)$ edges. However, if it so happens that the existing partial coloring of the graph only involves $O(\Delta)$ color types instead, then the same algorithm will only take $O(n\Delta)=O(m)$ time (by the near-regularity assumption) and we will be done. As such, the underlying idea behind our algorithm is to modify the current partial edge coloring (without decreasing the number of uncolored edges) so that the total number of color types reduces from $O(\Delta^{2})$ to $O(\Delta)$ only.
To explore this direction, let us assume for now the input graph is bipartite, which greatly simplifies the structure of Vizing fans and Vizing chains, thus allowing us to convey the key ideas more clearly; later we highlight some of the key challenges that arise when dealing with general graphs. We shall note that it has been known since the 80s that one can $\Delta$-edge color bipartite graphs in $\tilde{O}(m)$ time [CH82, COS01]. However, the algorithms for bipartite graphs use techniques that are entirely different from Vizing fans and Vizing chains, and which do not involve solving the color extension problem at all. In particular, prior to this work, it was unclear whether one can efficiently solve the color extension problem in bipartite graphs. Therefore, the assumption of a bipartite input graph does not trivialize our goal of using Vizing fans and Vizing chains for efficiently solving the color extension problem. Additionally, we can assume that in the color extension problem, the last $O(m/\Delta)$ edges to be colored can be partitioned into $O(1)$ matchings (this guarantee follows immediately from the recursive framework we outlined earlier), and that we deal with each of these matchings separately. In other words, we can also assume that the uncolored edges are vertex-disjoint.
Let $\chi$ be a partial $(\Delta+1)$-edge coloring, and for any vertex $w\in V$, let $\mathsf{miss}_{\chi}(w)\subseteq[\Delta+1]$ be the set of colors missing from the edges incident to $w$ under $\chi$. Given any uncolored edge $(u,v)$, the color type of this edge $(u,v)$ would be $\{c_{u},c_{v}\}$ for some arbitrary choices of $c_{u}\in\mathsf{miss}_{\chi}(u)$ and $c_{v}\in\mathsf{miss}_{\chi}(v)$ (it is possible for an edge to be able to choose more than one color type, but we fix one arbitrary choice among them); in other words, if we flip the $\{c_{u},c_{v}\}$-alternating path starting at $u$, then we can assign $\chi(u,v)$ to be $c_{v}$. To reduce the total number of different color types to $O(\Delta)$, we would have to make some color types much more popular: at the beginning, a type spans an $\Omega(1/\Delta^{2})$ proportion of the uncolored edges but we would like to have a type spanning an $\Omega(1/\Delta)$ proportion. For this purpose, we fix an arbitrary color type $\{\alpha,\beta\}$, and want to modify $\chi$ to transform the type of an arbitrary uncolored edge $(u,v)$ from $\{c_{u},c_{v}\}$ to $\{\alpha,\beta\}$ – we call this popularizing the edge $(u,v)$. To do this, we can simply flip the $\{\alpha,c_{u}\}$-alternating path $P_{u}$ starting at $u$ and the $\{\beta,c_{v}\}$-alternating path $P_{v}$ starting at $v$.
There are two technical issues regarding this path-flipping approach. Firstly, the alternating paths $P_{u}$ and $P_{v}$ could be very long, and require a long time for being flipped. More importantly, flipping $P_{u}$ and $P_{v}$ could possibly damage other $\{\alpha,\beta\}$-type (uncolored) edges that we popularized before. More specifically, say that we have popularized a set $\Phi$ of uncolored edges. However, when popularizing the next uncolored edge $(u,v)$, it could be the case that the $\{\alpha,c_{u}\}$-alternating path $P_{u}$ is ending at a vertex $u^{\prime}$ for some edge $(u^{\prime},v^{\prime})\in\Phi$. If we flip the path $P_{u}$, then $(u^{\prime},v^{\prime})$ would no longer be of $\{\alpha,\beta\}$-type as $\alpha$ would not be missing at $u^{\prime}$ anymore.
Our key observation is that when $|\Phi|$ is relatively small, most choices for an alternating path $P_{u}$ cannot be ending at edges in $\Phi$. Consider the above bad example where $P_{u}$ is a $\{c_{u},\alpha\}$-alternating path ending at $u^{\prime}$ for some $(u^{\prime},v^{\prime})\in\Phi$. Let us instead look at this from the perspective of the $\{\alpha,\beta\}$-type edge $(u^{\prime},v^{\prime})\in\Phi$. For any $(u^{\prime},v^{\prime})\in\Phi$ and any color $\gamma$, there can be at most one uncolored edge $(u,v)\notin\Phi$ whose corresponding path $P_{u}$ is the same $\{\gamma,\alpha\}$-alternating path starting at $u^{\prime}$; this is because any vertex belongs to at most one alternating path of a fixed type (here, the type $\{\gamma,\alpha\}$) and each vertex belongs to at most one uncolored edge (recall that the uncolored edges form a matching). This is also true for $\{\gamma,\beta\}$-type edges for the same exact reason. As such, ranging over all possible choices of $\gamma\in[\Delta+1]$, there are at most $O(|\Phi|\Delta)$ uncolored edges $(u,v)$ whose alternating paths could damage the set $\Phi$. Therefore, as long as the size of $\Phi$ is a $o(1/\Delta)$ fraction (or more precisely, a $O(1/\Delta)$ fraction, for a sufficiently small constant hiding in the $O$-notation) of all uncolored edges, the following property holds: a constant fraction of uncolored edges $e\notin\Phi$ can be popularized using the above method without damaging $\Phi$.
This property resolves both technical issues raised earlier simultaneously. Let $\lambda$ denote the number of uncolored edges. For the second issue, as long as $|{\Phi}|=o(\lambda/\Delta)$, we can take a random uncolored edge $(u,v)\notin\Phi$ and flip $P_{u}$ and $P_{v}$ if this does not damage any already popularized edge in $\Phi$; by the observation above, this happens with constant probability. For the first issue, we can show that the expected length of alternating paths $P_{u}$ and $P_{v}$, for the random uncolored edge picked above, is $O(m/\lambda)$; indeed, this is because the number of edges colored $\alpha$ or $\beta$ is $O(n)$, hence the total length of all alternating paths with one color being fixed to $\alpha$ or $\beta$ is $O(m)$. All in all, the total time we spend to popularize a single type $\{\alpha,\beta\}$ to become a $\Theta(1/\Delta)$ fraction of all uncolored edges is $O(\lambda/\Delta\cdot m/\lambda)=O(m/\Delta)$. Since coloring edges of a single type takes $O(n)=O(m/\Delta)$ time by our earlier discussion, we can color in this way a $\Theta(1/\Delta)$ fraction of all uncolored edges in $O(m/\Delta)$ expected time. As a direct corollary, we can color all uncolored edges in $O(m\log{n})$ time, hence solving the color extension problem, in (near-regular) bipartite graphs, in near-linear time. The above argument will be detailed in the proof of Lemma 4.2 in Section 6.
Collecting U-Fans in General Graphs: We now discuss the generalization of our scheme above to non-bipartite graphs. Existence of odd cycles in non-bipartite graphs implies that we can no longer assign a color type $\{c_{u},c_{v}\}$ to an uncolored edge $(u,v)$ for $c_{u}\in\mathsf{miss}_{\chi}(u),c_{v}\in\mathsf{miss}_{\chi}(v)$, and hope that flipping the $\{c_{u},c_{v}\}$-alternating path from $u$ allows us to color the edge $(u,v)$ with $c_{v}$ (because the path may end at $v$, and thus after flipping it $c_{v}$ will no longer be missing from $v$). This is where Vizing fans and Vizing chains come into play: in non-bipartite graphs, a color type of an uncolored edge $(u,v)$ is $\{c_{u},\gamma_{u,v}\}$ where $c_{u}$ is an arbitrary color in $\in\mathsf{miss}_{\chi}(u)$ but $\gamma_{u,v}$ is determined by the Vizing fan around $u$ and the alternating path that we take for coloring $(u,v)$ (or the Vizing chain of $(u,v)$). Thus, while as before we can switch $c_{u}$ with some fixed color $\alpha$, it is unclear how to flip alternating paths to change $\gamma_{u,v}$ to $\beta$ also, in order to popularize the edge $(u,v)$ to be of some designated type $\{\alpha,\beta\}$ —without damaging another popularized edge as a result.
To address this challenge, we rely on the notion of a u-fan, introduced by [GNK+85], which is the non-bipartite graph analogue of an uncolored edge in bipartite graphs. A u-fan of type $\{\alpha,\beta\}$ is a pair of uncolored edges $(u,v),(u,w)$ such that $\alpha\in\mathsf{miss}_{\chi}(u)$ and $\beta\in\mathsf{miss}_{\chi}(u)\cap\mathsf{miss}_{\chi}(v)$. Consider the $\{\alpha,\beta\}$-alternating path starting at $u$. As at least one of $v$ or $w$ is not the other endpoint of this alternating path (say $v$), flipping this path allows us to assign the color $\beta$ to $(u,v)$. Consequently, as u-fans are similar to edges in bipartite graphs, we can still essentially (but not exactly) apply our color type reduction approach if all the uncolored edges are paired as u-fans. Therefore, it suffices to modify $\chi$ to pair all uncolored edges together to form u-fans.
In order to pair different uncolored edges together and form $u$-fans, we should first be able to move uncolored edges around. Such operations already appeared in some previous work on dynamic edge coloring [DHZ19, Chr23, Chr24]. Basically, for any uncolored edge $(u,v)$, we can modify $\chi$ to shift this uncolored edge to any position on its Vizing chain. This naturally leads to a win-win situation: if the Vizing chain is short, then $(u,v)$ can be colored efficiently using Vizing’s procedure; otherwise if most Vizing chains are long, then there must be a pair of Vizing chains meeting together after a few steps, so we can shift two uncolored edges to form a u-fan efficiently.
Let us make this a bit more precise. Fix an arbitrary color $\alpha$ and consider the set $U_{\alpha}$ of all the uncolored edges $(u,v)$ such that $\alpha\in\mathsf{miss}_{\chi}(u)$ and the respective Vizing chain is of type $\{\alpha,\cdot\}$. Also, assume there are $m_{\alpha}$ edges colored $\alpha$ under $\chi$. If most $\{\alpha,\cdot\}$-Vizing chains have length larger than $\Omega(m_{\alpha}/|U_{\alpha}|)$, then on average, two Vizing chains should meet within the first $O(m_{\alpha}/|U_{\alpha}|)$ steps; in this case, we can repeatedly pick two intersecting Vizing chains and create a u-fan by shifting their initial uncolored edges to the intersection of these Vizing chains. Given the length of the chains, this takes $O(m_{\alpha}/|U_{\alpha}|)$ time. Otherwise, the average cost of applying Vizing’s color extension procedure is $O(m_{\alpha}/|U_{\alpha}|)$, and in this case we can directly color all those edges in $O(m_{\alpha})$ time. Summing over all different $\alpha\in[\Delta+1]$ gives a near-linear runtime. The above argument will be detailed in the proof of Lemma 4.1 in Section 5.
The above discussion leaves out various technical challenges. For instance, moving around uncolored edges as described above breaks the assumption that the uncolored edges form a matching. Handling this requires dedicating different colors from $\mathsf{miss}_{\chi}(u)$ for every uncolored edge incident on a vertex $u$. This is formalized via the notion of separability in Section 3.1. Additionally, we have ignored all algorithmic aspects of (efficiently) finding pairs of intersecting Vizing chains, as well as the corner cases of Vizing fan intersections and fan-chain intersections. We defer the discussions on these details to the main body of the paper.
2 Preliminaries
We begin by describing the basic notation used throughout the paper in Section 2.1 and then give the classic definitions of Vizing fans and Vizing chains in Section 2.2. In Section 3, we introduce the new definitions that provide the basic building blocks for our algorithm.
2.1 Notation
Let $G=(V,E)$ be graph on $n$ vertices with $m$ edges and maximum degree $\Delta$ and let $\chi:E\longrightarrow[\Delta+1]\cup\{\bot\}$ be a (partial) $(\Delta+1)$-edge coloring of $G$. We refer to edges $e\in E$ with $\chi(e)=\bot$ as uncolored. Given a vertex $u\in V$, we denote the set of colors that are not assigned to any edge incident on $u$ by $\mathsf{miss}_{\chi}(u)$. We sometimes refer to $\mathsf{miss}_{\chi}(u)$ as the palette of $u$. We say that the colors in $\mathsf{miss}_{\chi}(u)$ are missing (or available) at $u$.
Given a path $P=e_{1},\dots,e_{k}$ in $G$, we say that $P$ is an $\{\alpha,\beta\}$-alternating path if $\chi(e_{i})=\alpha$ whenever $i$ is odd and $\chi(e_{i})=\beta$ whenever $i$ is even (or vice versa). We say that the alternating path $P$ is maximal if one of the colors $\alpha$ or $\beta$ is missing at each of the endpoints of $P$. We refer to the process of changing the color of each edge $e_{i}\in P$ with color $\alpha$ (resp. $\beta$) to $\beta$ (resp. $\alpha$) as flipping the path $P$. We denote by $|P|$ the length (i.e., the number of edges) of the alternating path $P$. We define the length $i$ prefix of the path $P$ to be the path $P_{\leq i}:=e_{1},\dots,e_{i}$.
Given an uncolored edge $e$ under $\chi$, by extending $\chi$ to $e$ we mean modifying $\chi$ so that $\chi(e)\neq\bot$ without creating new uncolored edges. Our algorithm will always work by modifying a partial coloring $\chi$; unless explicitly specified otherwise, every new concept we define (such as u-fans and separable collection) will be defined with respect to this particular partial coloring $\chi$.
2.2 Vizing fans and chains
Vizing Fans: We now define the notion of Vizing fans, which has been used extensively in the edge coloring literature [Viz64, GNK+85, Sin19].
Definition 2.1 (Vizing fan).
A Vizing fan is a sequence $\operatorname*{\mathbf{F}}=(u,\alpha),(v_{1},c_{1}),\dots,(v_{k},c_{k})$ where $u,v_{1},\dots,v_{k}$ are distinct vertices and $c_{1},\dots,c_{k}$ are colors such that
  1. 1.
    $\alpha\in\mathsf{miss}_{\chi}(u)$ and $c_{i}\in\mathsf{miss}_{\chi}(v_{i})$ for all $i\in[k]$.
  2. 2.
    $v_{1},\dots,v_{k}$ are distinct neighbours of $u$.
  3. 3.
    $\chi(u,v_{1})=\bot$ and $\chi(u,v_{i})=c_{i}$ for all $i>1$.
  4. 4.
    Either $c_{k}\in\mathsf{miss}_{\chi}(u)$ or $c_{k}\in\{c_{1},\dots,c_{k-1}\}$.
We say that the Vizing fan $\operatorname*{\mathbf{F}}=(u,\alpha),(v_{0},c_{0}),\dots,(v_{k},c_{k})$ is $\alpha$-primed, has center $u$ and leaves $v_{1},\dots,v_{k}$. We refer to $c_{i}$ as the color of $v_{i}$ within $\operatorname*{\mathbf{F}}$. A crucial property is that we can rotate colors around the Vizing fan $\operatorname*{\mathbf{F}}$ by setting $\chi(u,v_{1})\leftarrow c_{1},\dots,\chi(u,v_{i-1})\leftarrow c_{i-1},\chi(u,v_{i})\leftarrow\bot$ for any $i\in[k]$. We say that $\operatorname*{\mathbf{F}}$ is a trivial Vizing fan if $c_{k}\in\mathsf{miss}_{\chi}(u)$. Note that, if $\operatorname*{\mathbf{F}}$ is trivial, we can immediately extend the coloring $\chi$ to $(u,v_{1})$ by rotating all the colors around $\operatorname*{\mathbf{F}}$ and setting $\chi(u,v_{k})\leftarrow c_{k}$.
Algorithm 1 describes the standard procedure used to construct Vizing fans. As input, it takes a vertex $u$ and a color $\alpha\in\mathsf{miss}_{\chi}(u)$, and returns an $\alpha$-primed Vizing fan with center $u$.
1 For each $x\in V$, let $\mathsf{clr}(x)\in\mathsf{miss}_{\chi}(x)$
2 $k\leftarrow 1$ and $v_{1}\leftarrow v$
3 $c_{1}\leftarrow\mathsf{clr}(v_{1})$
4 while $c_{k}\notin\{c_{1},\dots,c_{k-1}\}$ and $c_{k}\notin\mathsf{miss}_{\chi}(u)$ do
5       Let $(u,v_{k+1})$ be the edge with color $\chi(u,v_{k+1})=c_{k}$
6       $c_{k+1}\leftarrow\mathsf{clr}(v_{k+1})$
7       $k\leftarrow k+1$
8
return $(u,\alpha),(v_{1},c_{1}),\dots,(v_{k},c_{k})$
[…]
The Algorithm Vizing: We now describe the algorithm Vizing that, given a Vizing fan $\operatorname*{\mathbf{F}}=(u,\alpha),(v_{1},c_{1}),\dots,(v_{k},c_{k})$ as input, extends the coloring $\chi$ to the edge $(u,v_{1})$ by building a Vizing chain. Algorithm 2 gives a formal description of this procedure.
1 if $\operatorname*{\mathbf{F}}$ is trivial then
2       $\chi(u,v_{i})\leftarrow c_{i}$ for all $i\in[k]$
3       return
4
5Let $P$ denote the maximal $\{\alpha,c_{k}\}$-alternating path starting at $u$
6 Extend $\chi$ to $(u,v_{1})$ by flipping the path $P$ and rotating colors in $\operatorname*{\mathbf{F}}$ (details in Lemma 2.2)
Algorithm 2 Vizing$(\operatorname*{\mathbf{F}})$
Thus, running $\textnormal{{Vizing}}(\operatorname*{\mathbf{F}})$ extends the coloring $\chi$ to the uncolored edge $(u,v_{1})$ by rotating colors in the Vizing fan $\operatorname*{\mathbf{F}}$ and flipping the colors of the alternating path $P$. We sometimes refer to the process of running $\textnormal{{Vizing}}(\operatorname*{\mathbf{F}})$ as activating the Vizing fan $\operatorname*{\mathbf{F}}$.
Lemma 2.2.
Algorithm 2 extends the coloring $\chi$ to the edge $(u,v_{1})$ in time $O(\Delta+|P|)$.
Proof.
To see that the path $P$ is well-defined, note that $\alpha\in\mathsf{miss}_{\chi}(u)$ and $c_{k}\notin\mathsf{miss}_{\chi}(u)$, so there is an $\{\alpha,c_{k}\}$-alternating path starting at $u$.
Extending the coloring $\chi$: By the definition of a Vizing fan, $c_{k}$ is the first repetition of some color in $\{c_{1},\ldots,c_{k-1}\}$, so there is a unique index $j\in[k-1]$ such that $c_{j}=c_{k}$, and $P$ has $(u,v_{j+1})$ as its first edge with color $c_{j}$. We consider the cases where the path $P$ does or does not have $v_{j}$ as an endpoint. If $P$ does not end at $v_{j}$, then we can rotate the colors of the first $j+1$ edges in the fan by setting $\chi(u,v_{1})\leftarrow c_{1},\dots,\chi(u,v_{j})\leftarrow c_{j}$ and flip the colors of the alternating path $P$. If $P$ does end at $v_{j}$, then we flip the colors of the alternating path $P$, rotate the colors of the fan by setting $\chi(u,v_{1})\leftarrow\chi(u,v_{2}),\dots,\chi(u,v_{k-1})\leftarrow\chi(u,v_{k})$, and set $\chi(u,v_{k})\leftarrow c_{k}$. Note that, while rotating the fan in this last step, we have $\chi(u,v_{j+1})=\alpha\neq c_{k}$.
Using standard data structures (see Section 7), we can implement this algorithm to run in time proportional to the number of edges that change their colors, which is $O(\Delta+|P|)$. ∎
[…]
3 Basic Building Blocks
In this section, we introduce the notation of u-fans, u-edges and separable collections, which are the definitions that work as the basic building blocks for our algorithms.
3.1 U-Fans and U-Edges
We begin by defining the notion of u-fans that was used by [GNK+85]. [^2] The term ‘u-fan’ was originally introduced by [GNK+85] as an abbreviation for ‘uncolored fan’.
Definition 3.1 (u-fan, [GNK+85]).
A u-fan is a tuple $\operatorname*{\boldsymbol{f}}=(u,v,w,\alpha,\beta)$ where $u$, $v$ and $w$ are distinct vertices and $\alpha$ and $\beta$ are distinct colors such that:
  1. 1.
    $(u,v)$ and $(u,w)$ are uncolored edges.
  2. 2.
    $\alpha\in\mathsf{miss}_{\chi}(u)$ and $\beta\in\mathsf{miss}_{\chi}(v)\cap\mathsf{miss}_{\chi}(w)$.
We say that $u$ is the center of $\operatorname*{\boldsymbol{f}}$ and that $v$ and $w$ are the leaves of $\operatorname*{\boldsymbol{f}}$. We also say that the u-fan $\operatorname*{\boldsymbol{f}}$ is $\{\alpha,\beta\}$-primed and define $c_{\operatorname*{\boldsymbol{f}}}(u):=\alpha$, $c_{\operatorname*{\boldsymbol{f}}}(v):=\beta$ and $c_{\operatorname*{\boldsymbol{f}}}(w):=\beta$ (i.e. given a vertex $x\in\operatorname*{\boldsymbol{f}}$, $c_{\operatorname*{\boldsymbol{f}}}(x)$ is the available color that $\operatorname*{\boldsymbol{f}}$ ‘assigns’ to $x$).
Activating U-Fans: Let $\operatorname*{\boldsymbol{f}}$ be an $\{\alpha,\beta\}$-primed u-fan with center $u$ and leaves $v$ and $w$. The key property of u-fans is that at most one of the $\{\alpha,\beta\}$-alternating paths starting at $v$ or $w$ ends at $u$. Say that the $\{\alpha,\beta\}$-alternating path starting at $v$ does not end at $u$. Then, after flipping this $\{\alpha,\beta\}$-alternating path, both $u,v$ are missing color $\alpha$. Thus, we can extend the coloring $\chi$ by assigning $\chi(u,v)\leftarrow\alpha$. We refer to this as activating the u-fan $\operatorname*{\boldsymbol{f}}$.
We also define the notion of a u-edge similarly to u-fans.
Definition 3.2 (u-edge).
A u-edge is a tuple $\operatorname*{\boldsymbol{e}}=(u,v,\alpha)$ where $(u,v)$ is an uncolored edge and $\alpha$ is a color such that $\alpha\in\mathsf{miss}_{\chi}(u)$.
We say that $u$ is the center of $\operatorname*{\boldsymbol{e}}$ and that $\alpha$ is the active color of $\operatorname*{\boldsymbol{e}}$. For notational convenience, we also say that the u-edge $\operatorname*{\boldsymbol{e}}$ is $\alpha$-primed and define $c_{\operatorname*{\boldsymbol{e}}}(u):=\alpha$ and $c_{\operatorname*{\boldsymbol{e}}}(v)=\bot$. [^3] Whenever we refer to an ‘uncolored edge $e$’, we are referring to an edge $e\in E$ such that $\chi(e)=\bot$, whereas a ‘u-edge $\operatorname*{\boldsymbol{e}}$’ always refers to the object from LABEL:{def:u-edge} and is denoted in bold.
Collections of U-Components: While working with both u-fans and u-edges simultaneously, we sometimes refer to some $\operatorname*{\boldsymbol{g}}$ that is either a u-fan or a u-edge as a u-component. Throughout this paper, we often consider collections of u-components $\operatorname*{\mathcal{U}}$. We only use the term ‘collection’ in this context, so we abbreviate ‘collection of u-components’ by just ‘collection’. We will be particularly interested in collections satisfying the following useful property, which we refer to as separability.
Definition 3.3 (Separable Collection).
Let $\chi$ be a partial $(\Delta+1)$-edge coloring of $G$ and $\mathcal{U}$ be a collection of u-components (i.e. u-fans and u-edges). We say that the collection $\mathcal{U}$ is separable if the following holds:
  1. 1.
    All of the u-components in $\mathcal{U}$ are edge-disjoint.
  2. 2.
    For each $x\in V$, the colors in the multi-set $C_{\mathcal{U}}(x):=\{c_{\operatorname*{\boldsymbol{g}}}(x)\mid\operatorname*{\boldsymbol{g}}\in\mathcal{U},\,x\in\operatorname*{\boldsymbol{g}}\}$ are distinct.
Damaged U-Components: Suppose we have a partial $(\Delta+1)$-edge coloring $\chi$ and a separable collection $\operatorname*{\mathcal{U}}$ w.r.t. $\chi$. Now, suppose that we modify the coloring $\chi$. We say that a u-component $\operatorname*{\boldsymbol{g}}\in\operatorname*{\mathcal{U}}$ is damaged if it is no longer a u-component w.r.t. the new coloring $\chi$.
We note that this can only happen for one of the following two reasons: (1) one of the uncolored edges in $\operatorname*{\boldsymbol{g}}$ is now colored, or (2) there is a vertex $x\in\operatorname*{\boldsymbol{g}}$ such the color $c_{\operatorname*{\boldsymbol{g}}}(x)$ that $\operatorname*{\boldsymbol{g}}$ assigns to $x$ is no longer missing at $x$.
The following lemma shows that flipping the colors of an alternating path cannot damage many u-components in a separable collection $\operatorname*{\mathcal{U}}$.
Lemma 3.4.
Let $\chi$ be a partial $(\Delta+1)$-edge coloring of a graph $G$, $\operatorname*{\mathcal{U}}$ a separable collection and $P$ a maximal alternating path in $\chi$. Then flipping the colors of the alternating path $P$ damages at most $2$ u-components in $\operatorname*{\mathcal{U}}$ (corresponding to the two endpoints of the path).
Proof.
Let $x$ and $y$ be the endpoints of the path $P$. Since only the palettes of $x$ and $y$ change after flipping the colors of $P$, only u-components $\operatorname*{\boldsymbol{g}}\in\operatorname*{\mathcal{U}}$ that contain $x$ and $y$ may be damaged. Since the palette of $x$ changes by $1$ and the collection $\operatorname*{\mathcal{U}}$ is separable, only one $\operatorname*{\boldsymbol{g}}\in\operatorname*{\mathcal{U}}$ containing $x$ may be damaged by the color $c_{\operatorname*{\boldsymbol{g}}}(x)$ no longer being available at $x$. The same holds for the vertex $y$. Thus, at most 2 u-components in $\operatorname*{\mathcal{U}}$ are damaged by this operation. ∎
3.2 Data Structures
In Section 7, we describe the data structures that we use to implement our algorithm. On top of the standard data structures used to maintain the edge coloring $\chi$, which allows us to implement Algorithms 1 and 2 efficiently, we also use data structures that allow us to efficiently maintain a separable collection $\operatorname*{\mathcal{U}}$. More specifically, the data structures that we use to implement a separable collection $\operatorname*{\mathcal{U}}$ support the following queries.
  • $\textsc{Insert}_{\operatorname*{\mathcal{U}}}(\operatorname*{\boldsymbol{g}})$: The input to this query is a u-component $\operatorname*{\boldsymbol{g}}$. In response, the data structure adds $\operatorname*{\boldsymbol{g}}$ to $\operatorname*{\mathcal{U}}$ if $\operatorname*{\mathcal{U}}\cup\{\operatorname*{\boldsymbol{g}}\}$ is separable and outputs fail otherwise.
  • $\textsc{Delete}_{\operatorname*{\mathcal{U}}}(\operatorname*{\boldsymbol{g}})$: The input to this query is a u-component $\operatorname*{\boldsymbol{g}}$. In response, the data structure removes $\operatorname*{\boldsymbol{g}}$ from $\operatorname*{\mathcal{U}}$ if $\operatorname*{\boldsymbol{g}}\in\operatorname*{\mathcal{U}}$ and outputs fail otherwise.
[…]
The following claim shows that it is always possible to answer a Missing-Color query.
Claim 3.5.
For each $x\in V$, the set $\mathsf{miss}_{\chi}(x)\setminus C_{\operatorname*{\mathcal{U}}}(x)$ is non-empty.
Proof.
Let $d$ denote the number of uncolored edges incident on $x$. Since the collection $\operatorname*{\mathcal{U}}$ is separable, we have that $|C_{\operatorname*{\mathcal{U}}}(x)|\leq d$. Since $|\mathsf{miss}_{\chi}(x)|\geq d+1$, it follows that $|C_{\operatorname*{\mathcal{U}}}(x)|<|\mathsf{miss}_{\chi}(x)|$ and so $\mathsf{miss}_{\chi}(x)\setminus C_{\operatorname*{\mathcal{U}}}(x)\neq\varnothing$. ∎
Furthermore, the data structure supports the following initialization operation.
  • $\textsc{Initialize}(G,\chi)$: Given a graph $G$ and an edge coloring $\chi$ of $G$, we can initialize the data structure with an empty separable collection $\operatorname*{\mathcal{U}}=\varnothing$.
In Section 7, we show how to implement the initialization operation in $O(m)$ time and each of these queries in $O(1)$ time with the appropriate data structures. These queries provide the ‘interface’ via which our algorithm will interact with the u-components.
A Remark on Randomization: In order to get good space complexity and running time simultaneously, we implement the standard data structures (used for Algorithms 1 and 2) and the data structure supporting the preceding queries using hashmaps (see Section 7). The following proposition describes the hashmaps used by our data structures. [^5] We could have even used the construction of [Kus22] to obtain exponentially small error probability in this case, but since we do not need this stronger guarantee we stick with the simpler work of [DadH90].
Proposition 3.6 ([DadH90]).
There exists a dynamic dictionary that, given a parameter $k$, can handle $k$ insertions, deletions, and map operations, uses $O(k)$ space and takes $O(1)$ worst-case time per operation with probability at least $1-O(1/k^{7})$.
This implementation gives us the guarantee that across the run of the entire algorithm, every query made to one of these data structures takes $O(1)$ time and each initialization operation takes $O(m)$ time, with high probability. Since the randomization used for the hashmaps is independent of the randomization used in the rest of the algorithm, we implicitly condition on the high probability event that, every operation performed using a hashmap runs in $O(1)$ time throughout the rest of the paper. [^6] Alternatively, one can replace these hashmaps with balanced search trees to make these data structures deterministic, incurring $O(\log n)$ overhead in the running time of each operation.
3.3 Vizing Fans in Separable Collections
Within our algorithm, we only construct Vizing fans and Vizing chains in a setting where there is some underlying separable collection $\operatorname*{\mathcal{U}}$. To ensure that activating and rotating colors around Vizing fans does not damage too many u-components, we choose the missing colors involved in Vizing fan constructions so that they ‘avoid’ the colors assigned to the u-components in $\operatorname*{\mathcal{U}}$.
Definition 3.7.
Let $\operatorname*{\mathcal{U}}$ be a separable collection and $\operatorname*{\mathbf{F}}=(u,\alpha),(v_{1},c_{1}),\dots,(v_{k},c_{k})$ be a Vizing fan. We say that the Vizing fan $\operatorname*{\mathbf{F}}$ if $\operatorname*{\mathcal{U}}$-avoiding if $c_{i}\in\mathsf{miss}_{\chi}(v_{i})\setminus C_{\operatorname*{\mathcal{U}}}(v_{i})$ for each leaf $v_{i}\in\operatorname*{\mathbf{F}}$.
We say that a Vizing fan $\operatorname*{\mathbf{F}}$ is a Vizing fan of the u-edge $\operatorname*{\boldsymbol{e}}=(u,v,\alpha)$ if $\operatorname*{\mathbf{F}}$ is $\alpha$-primed, has center $u$ and its first leaf is $v$. The following lemma shows that we can always find a $\operatorname*{\mathcal{U}}$-avoiding Vizing fan for a u-edge.
Lemma 3.8.
Given a u-edge $\operatorname*{\boldsymbol{e}}\in\operatorname*{\mathcal{U}}$, there exists a $\operatorname*{\mathcal{U}}$-avoiding Vizing fan $\operatorname*{\mathbf{F}}$ of $\operatorname*{\boldsymbol{e}}$. Furthermore, we can compute such a Vizing fan in $O(\Delta)$ time.
Proof.
By 3.5, we can always find a collection of colors $\{\mathsf{clr}(x)\}_{x\in V}$ such that $\mathsf{clr}(x)\in\mathsf{miss}_{\chi}(x)\setminus C_{\operatorname*{\mathcal{U}}}(x)$ for each $x\in V$. If we construct a Vizing fan $\operatorname*{\mathbf{F}}$ by calling Vizing-Fan and using such a collection of colors within the algorithm, then it follows that the Vizing fan $\operatorname*{\mathbf{F}}$ is $\operatorname*{\mathcal{U}}$-avoiding. By combining the standard implementation of Algorithm 1 with the queries described in Section 3.2 we can do this in $O(\Delta)$ time. ∎
The following lemma describes some crucial properties of $\operatorname*{\mathcal{U}}$-avoiding Vizing fans.
Lemma 3.9.
Let $\chi$ be a $(\Delta+1)$-edge coloring of a graph $G$ and $\operatorname*{\mathcal{U}}$ be a separable collection. For any u-edge $\operatorname*{\boldsymbol{e}}=(u,v,\alpha)\in\operatorname*{\mathcal{U}}$ with a $\operatorname*{\mathcal{U}}$-avoiding Vizing fan $\operatorname*{\mathbf{F}}$, we have the following:
  1. 1.
    Rotating colors around $\operatorname*{\mathbf{F}}$ does not damage any u-component in $\operatorname*{\mathcal{U}}\setminus\{\operatorname*{\boldsymbol{e}}\}$.
  2. 2.
    Calling $\textnormal{{Vizing}}(\operatorname*{\mathbf{F}})$ damages at most one u-component in $\operatorname*{\mathcal{U}}\setminus\{\operatorname*{\boldsymbol{e}}\}$. Furthermore, we can identify the damaged u-component in $O(1)$ time.
Proof.
Let $\operatorname*{\mathbf{F}}=(u,\alpha),(v_{1},c_{1}),\dots,(v_{k},c_{k})$. Rotating colors around $\operatorname*{\mathbf{F}}$ will only remove the color $c_{i}$ from the palette of a leaf $v_{i}$ appearing in $\operatorname*{\mathbf{F}}$. Since $c_{i}\notin C_{\operatorname*{\mathcal{U}}}(v_{i})$, these changes to the palettes will not damage any u-component in $\operatorname*{\mathcal{U}}$. However, rotating colors around $\operatorname*{\mathbf{F}}$ will color the edge $(u,v)$, damaging the u-edge $\operatorname*{\boldsymbol{e}}$.
Now, suppose we call $\textnormal{{Vizing}}(\operatorname*{\mathbf{F}})$. If the Vizing fan $\operatorname*{\mathbf{F}}$ is trivial, then the algorithm rotates all the colors in $\operatorname*{\mathbf{F}}$ and sets $\chi(u,v_{k})\leftarrow c_{k}$. Rotating the colors can only damage $\operatorname*{\boldsymbol{e}}\in\operatorname*{\mathcal{U}}$, and removing $c_{k}$ from the palette of $u$ can damage at most one other u-component in $\operatorname*{\mathcal{U}}$ since $\operatorname*{\mathcal{U}}$ is separable. Using the queries from Section 3.2, we can check if such a u-component exist and return it in $O(1)$ time. If $\operatorname*{\mathbf{F}}$ is not trivial, then the algorithm flips the $\{\alpha,c_{k}\}$-alternating path $P=\textnormal{{Vizing-Path}}(\operatorname*{\mathbf{F}})$ and rotates colors around $\operatorname*{\mathbf{F}}$. By similar arguments, rotating colors around $\operatorname*{\mathbf{F}}$ can only damage $\operatorname*{\boldsymbol{e}}\in\operatorname*{\mathcal{U}}$. Let $x$ denote the endpoint of $P$ that is not $u$. Flipping the path $P$ might remove either the color $\alpha$ or $c_{k}$ from the palette of $x$, which might damage at most one u-component in $\operatorname*{\mathcal{U}}\setminus\{e\}$ since $\operatorname*{\mathcal{U}}$ is separable. We can again identify this u-component in $O(1)$ time using the queries from Section 3.2. It also removes the color $\alpha$ from the palette of $u$, but this cannot damage any u-component other than $\operatorname*{\boldsymbol{e}}$ (again, since $\operatorname*{\mathcal{U}}$ is separable). ∎
4 The Main Algorithm
Our final algorithm consists of two main components. The first component is an algorithm called Construct-U-Fans that takes a partial $(\Delta+1)$-edge coloring $\chi$ with $\lambda$ uncolored edges and either extends $\chi$ to $\Omega(\lambda)$ of these edges or modifies the coloring to construct a separable collection of $\Omega(\lambda)$ u-fans. Lemma 4.1 summarizes the behaviour of this algorithm.
Lemma 4.1.
Given a graph $G$, a partial $(\Delta+1)$-edge coloring $\chi$ of $G$ and a set of $\lambda$ uncolored edges $U$, the algorithm Construct-U-Fans does one of the following in $O((m+\Delta\lambda)\log n)$ time:
  1. 1.
    Extends the coloring to $\Omega(\lambda)$ uncolored edges.
  2. 2.
    Modifies $\chi$ to obtain a separable collection of $\Omega(\lambda)$ u-fans $\mathcal{U}$.
The second component is an algorithm called Color-U-Fans that takes a collection of $\lambda$ u-fans and extends the coloring to $\Omega(\lambda)$ of the edges in the u-fans. Lemma 4.2 summarizes the behaviour of this algorithm.
Lemma 4.2.
Given a graph $G$, a partial $(\Delta+1)$-edge coloring $\chi$ of $G$ and a separable collection of $\lambda$ u-fans $\mathcal{U}$, the algorithm Color-U-Fans extends $\chi$ to $\Omega(\lambda)$ edges in $O(m\log n)$ time w.h.p.
In Sections 5 and 6, we describe these algorithms and prove Lemmas 4.1 and 4.2 respectively. Using these lemmas, we now show how to efficiently extend an edge coloring $\chi$ to the remaining uncolored edges.
Lemma 4.3.
Given a graph $G$ and a partial $(\Delta+1)$-edge coloring $\chi$ of $G$ with $\lambda$ uncolored edges $U$, we can extend $\chi$ to the remaining uncolored edges in time $O((m+\Delta\lambda)\log^{2}n)$ w.h.p.
Proof.
Let $U$ denote the set of edges that are uncolored by $\chi$. We can then apply Lemma 4.1 to either extend $\chi$ to a constant fraction of the edges in $U$ or construct a separable collection of $\Omega(\lambda)$ u-fans $\mathcal{U}$ in $O((m+\Delta\lambda)\log n)$ time. In the second case, we can then apply Lemma 4.2 to color $\Omega(\lambda)$ of the edges contained in these u-fans in $O(m\log n)$ time w.h.p. Thus, we can extend $\chi$ to a constant fraction of the uncolored edges in $O((m+\Delta\lambda)\log n)$ time w.h.p. After repeating this process $O(\log\lambda)\leq O(\log n)$ many times, no edges remain uncolored. Thus, we can extend the coloring $\chi$ to the entire graph in $O((m+\Delta\lambda)\log^{2}n)$ time w.h.p. ∎
We now use this lemma to prove Theorem 1.1.
Proof of Theorem 1.1..
We prove this by applying Lemma 4.3 to the standard Euler partition framework [GNK+85, Sin19]. Given a graph $G$, we partition it into two edge-disjoint subgraphs $G_{1}$ and $G_{2}$ on the same vertex set such that $\Delta(G_{i})\leq\lceil\Delta/2\rceil$ for each $G_{i}$, where $\Delta(G_{i})$ denotes the maximum degree of $G_{i}$. We then recursively compute a $(\Delta(G_{i})+1)$-edge coloring $\chi_{i}$ for each $G_{i}$. Combining $\chi_{1}$ and $\chi_{2}$, we obtain a $(\Delta+3)$-edge coloring $\chi$ of $G$. We then uncolor the two smallest color classes in $\chi$, which contain $O(m/\Delta)$ edges, and apply Lemma 4.3 to recolor all of the uncolored edges in $\chi$ using only $\Delta+1$ colors in $O(m\log^{2}n)$ time w.h.p.
To show that the total running time of the algorithm is $O(m\log^{3}n)$, first note that the depth of the recursion tree is $O(\log\Delta)$. Next, consider the $i^{th}$ level of the recursion tree, for an arbitrary $i=O(\log\Delta)$: we have $2^{i}$ edge-disjoint subgraphs $G_{1},\dots,G_{2^{i}}$ such that $\Delta(G_{j})\leq O(\Delta/2^{i})$ and $\sum_{j=1}^{2^{i}}|E(G_{j})|=m$. Since the total running time at recursion level $i$ is $O(m\log^{2}n)$ and the depth of the recursion tree is $O(\log\Delta)$, it follows that the total running time is $O(m\log^{3}n)$ w.h.p. ∎
5 The Algorithm Construct-U-Fans (Proof of Lemma 4.1)
As input, the algorithm Construct-U-Fans is given a graph $G$ and a partial $(\Delta+1)$-edge coloring $\chi$ of $G$ with $\lambda$ uncolored edges. It begins by taking the $\lambda$ uncolored edges in $\chi$ and using them to construct a separable collection $\operatorname*{\mathcal{U}}$ of $\lambda$ u-edges in an obvious way, which we describe in Lemma 5.14. It then uses two subroutines, Prune-Vizing-Fans and Reduce-U-Edges, to reduce the number of u-edges in $\operatorname*{\mathcal{U}}$, either by coloring them or modifying $\chi$ to turn them into u-fans. More specifically, for each color $\alpha\in[\Delta+1]$, the algorithm considers the collection $\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$ of u-edges that are $\alpha$-primed and (1) calls the subroutine Prune-Vizing-Fans to ensure that the u-edges in $\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$ have vertex-disjoint Vizing fans by either coloring the u-edges with overlapping Vizing fans or using them to create u-fans, and (2) calls the subroutine Reduce-U-Edges which either extends the coloring to the u-edges in $\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$ or uses them to create u-fans. Algorithm 3 gives the pseudocode for this algorithm.
1 Construct a separable collection of $\lambda$ u-edges $\operatorname*{\mathcal{U}}$
2 for each $\alpha\in[\Delta+1]$ do
3       $\mathcal{F}\leftarrow\textnormal{{Prune-Vizing-Fans}}(\operatorname*{\mathcal{U}},\alpha)$
[…]
5
return $\operatorname*{\mathcal{U}}$
Algorithm 3 Construct-U-Fans
Organization of Section 5: We first describe and analyze the subroutines Prune-Vizing-Fans and Reduce-U-Edges used by the algorithm Construct-U-Fans before proving Lemma 4.1 in Section 5.3.
5.1 The Subroutine Prune-Vizing-Fans
As input, the subroutine Prune-Vizing-Fans is given a graph $G$, a partial $(\Delta+1)$-edge coloring $\chi$ of $G$, a color $\alpha$ and a separable collection $\operatorname*{\mathcal{U}}$ with $\lambda_{\alpha}$ $\alpha$-primed u-edges. We let $\operatorname*{\mathcal{E}}(\operatorname*{\mathcal{U}})\subseteq\operatorname*{\mathcal{U}}$ denote the set of u-edges in $\operatorname*{\mathcal{U}}$ and $\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})\subseteq\operatorname*{\mathcal{E}}(\operatorname*{\mathcal{U}})$ denote the set of u-edges in $\operatorname*{\mathcal{U}}$ that are $\alpha$-primed. The subroutine then modifies the coloring $\chi$ and collection $\operatorname*{\mathcal{U}}$ to ensure that the Vizing fans of the u-edges in $\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$ are vertex-disjoint.
Algorithm Description: The subroutine Prune-Vizing-Fans scans through the u-edges in $\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$ and constructs $\operatorname*{\mathcal{U}}$-avoiding Vizing fans at these u-edges (see Lemma 3.8), maintaining a vertex-disjoint subset $\mathcal{F}$ of these Vizing fans throughout the process. After constructing a Vizing fan $\operatorname*{\mathbf{F}}$ for a u-edge $\operatorname*{\boldsymbol{e}}\in\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$, the subroutine checks if $F$ intersects some other fan in $\mathcal{F}$. If not, it adds $\operatorname*{\mathbf{F}}$ to $\mathcal{F}$. Otherwise, it uses these intersecting Vizing fans to either extend the coloring $\chi$ or construct a u-fan, removing the corresponding Vizing fans and u-edges from $\mathcal{F}$ and $\operatorname*{\mathcal{U}}$ respectively.
More formally, let $\operatorname*{\boldsymbol{e}}_{1},\dots,\operatorname*{\boldsymbol{e}}_{\ell}$ denote the u-edges in $\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$ at the time when the subroutine is called and let $\operatorname*{\boldsymbol{e}}_{i}=(u_{i},v_{i},\alpha)$. Initialize an empty set of Vizing fans $\mathcal{F}\leftarrow\varnothing$. Then, for each $i\in[\ell]$, the subroutine constructs a $\operatorname*{\mathcal{U}}$-avoiding Vizing fan $\operatorname*{\mathbf{F}}_{i}$ of $\operatorname*{\boldsymbol{e}}_{i}$ and checks if any of the vertices in $\operatorname*{\mathbf{F}}_{i}$ (i.e. any of the leaves or the center of $\operatorname*{\mathbf{F}}_{i}$) are contained in any of the Vizing fans in $\mathcal{F}$. Note that, since $\operatorname*{\mathcal{U}}$ is separable and $u_{1},\dots,u_{\ell}$ are the centers of $\alpha$-primed u-edges, these vertices are all distinct. If $\operatorname*{\mathbf{F}}_{i}$ is vertex-disjoint from the Vizing fans in $\mathcal{F}$, we add $\operatorname*{\mathbf{F}}_{i}$ to $\mathcal{F}$. Otherwise, consider the following 3 cases.
  • If $u_{i}$ is a leaf in $\operatorname*{\mathbf{F}}_{j}$ for some $\operatorname*{\mathbf{F}}_{j}\in\mathcal{F}$: Then we rotate the fan $\operatorname*{\mathbf{F}}_{j}$ so that $(u_{i},u_{j})$ is uncolored, set $\chi(u_{i},u_{j})\leftarrow\alpha$, remove $\operatorname*{\mathbf{F}}_{j}$ from $\mathcal{F}$ and remove $\operatorname*{\boldsymbol{e}}_{i}$ and $\operatorname*{\boldsymbol{e}}_{j}$ from $\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$.
If this is not the case, then some leaf of $\operatorname*{\mathbf{F}}_{i}$ appears in a Vizing fan in $\mathcal{F}$. Let $w$ be the first such leaf in $\operatorname*{\mathbf{F}}_{i}$.
  • If $w=u_{j}$ for some $\operatorname*{\mathbf{F}}_{j}\in\mathcal{F}$: Then we rotate the fan $\operatorname*{\mathbf{F}}_{i}$ so that $(u_{i},w)$ is uncolored, set $\chi(u_{i},w)\leftarrow\alpha$, remove $\operatorname*{\mathbf{F}}_{j}$ from $\mathcal{F}$ and remove $\operatorname*{\boldsymbol{e}}_{i}$ and $\operatorname*{\boldsymbol{e}}_{j}$ from $\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$.
  • If $w$ is a leaf in $\operatorname*{\mathbf{F}}_{j}\in\mathcal{F}$: Then we rotate the fans $\operatorname*{\mathbf{F}}_{i}$ and $\operatorname*{\mathbf{F}}_{j}$ so that $(u_{i},w)$ and $(u_{j},w)$ are uncolored, create a u-fan $\operatorname*{\boldsymbol{f}}=(w,u_{i},u_{j},\beta,\alpha)$ for an arbitrary $\beta\in\mathsf{miss}_{\chi}(w)\setminus C_{\operatorname*{\mathcal{U}}}(w)$, remove $\operatorname*{\mathbf{F}}_{j}$ from $\mathcal{F}$, add $\operatorname*{\boldsymbol{f}}$ to $\operatorname*{\mathcal{U}}$ and remove $\operatorname*{\boldsymbol{e}}_{i}$ and $\operatorname*{\boldsymbol{e}}_{j}$ from $\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$. [^7] Note that, since the vertices $u_{1},\dots,u_{\ell}$ are all distinct, these cases are exhaustive.
The subroutine then returns the set $\mathcal{F}$ of vertex-disjoint Vizing fans.
Analysis of Prune-Vizing-Fans
The following lemmas summarize the main properties of the subroutine Prune-Vizing-Fans.
Lemma 5.1.
Prune-Vizing-Fans returns a set $\mathcal{F}$ of vertex-disjoint $\operatorname*{\mathcal{U}}$-avoiding Vizing fans for all of the u-edges in $\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$.
Proof.
We can show by induction that, after the subroutine finishes scanning the u-edge $\operatorname*{\boldsymbol{e}}_{i}$, the set $\mathcal{F}$ consists of vertex-disjoint $\operatorname*{\mathcal{U}}$-avoiding Vizing fans for the u-edges in $\{\operatorname*{\boldsymbol{e}}_{1},\dots,\operatorname*{\boldsymbol{e}}_{i}\}\cap\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$. [^8] Since we remove edges from $\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$ throughout this process, $\{\operatorname*{\boldsymbol{e}}_{1},\dots,\operatorname*{\boldsymbol{e}}_{i-1}\}$ might not be contained in $\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$. This is true trivially for $i=1$. Now, suppose that this is true for $i-1$. If the $\operatorname*{\mathcal{U}}$-avoiding Vizing fan $\operatorname*{\mathbf{F}}_{i}$ of $\operatorname*{\boldsymbol{e}}_{i}$ does not intersect any of the Vizing fans in $\mathcal{F}$, then we add $\operatorname*{\mathbf{F}}_{i}$ to $\mathcal{F}$ and are done.
Thus, we assume this is not the case. Let $x_{0}$ denote $u_{i}$ and $x_{1},\dots,x_{k}$ denote the leaves of the Vizing fan $\operatorname*{\mathbf{F}}_{i}$ (in order). Let $x_{p}$ be the first vertex in the sequence $x_{0},x_{1},\dots,x_{k}$ which appears in some Vizing fan $\operatorname*{\mathbf{F}}_{j}\in\mathcal{F}$. It follows from our algorithm that we remove $\operatorname*{\boldsymbol{e}}_{i}$ and $\operatorname*{\boldsymbol{e}}_{j}$ from $\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$ while only affecting the palettes of vertices in $\operatorname*{\mathbf{F}}_{j}$ and $x_{0},\dots,x_{p}$ by rotating the fans $\operatorname*{\mathbf{F}}_{i}$ and $\operatorname*{\mathbf{F}}_{j}$. Since all of the Vizing fans in $\mathcal{F}$ were vertex-disjoint, changing the palettes of vertices in $\operatorname*{\mathbf{F}}_{j}$ cannot affect the palettes of vertices in any of the Vizing fans in $\mathcal{F}\setminus\{\operatorname*{\mathbf{F}}_{j}\}$, and hence they are still valid. Similarly, none of the vertices $x_{0},\dots,x_{p}$ are contained in any of the Vizing fans in $\mathcal{F}\setminus\{\operatorname*{\mathbf{F}}_{j}\}$, so changing their palettes does not affect the validity of any other Vizing fans in $\mathcal{F}$. Thus, after removing $\operatorname*{\mathbf{F}}_{j}$ from $\mathcal{F}$, removing $\operatorname*{\boldsymbol{e}}_{i}$ and $\operatorname*{\boldsymbol{e}}_{j}$ from $\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$ and making the changes to $\chi$, the set $\mathcal{F}$ consists of vertex-disjoint Vizing fans of the u-edges in $\{\operatorname*{\boldsymbol{e}}_{1},\dots,\operatorname*{\boldsymbol{e}}_{i}\}\cap\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$. To see why all of these Vizing fans also remain $\operatorname*{\mathcal{U}}$-avoiding, note that any u-fan $\operatorname*{\boldsymbol{f}}$ added to $\operatorname*{\mathcal{U}}$ is vertex-disjoint from these Vizing fans by an analogous argument. ∎
Lemma 5.2.
The subroutine Prune-Vizing-Fans maintains the invariant that $\operatorname*{\mathcal{U}}$ is separable.
Proof.
We can show by induction that, after the subroutine finishes scanning the u-edge $\operatorname*{\boldsymbol{e}}_{i}$, the collection $\operatorname*{\mathcal{U}}$ is separable. This is true trivially for $i=1$, since the subroutine does not change $\chi$ or $\operatorname*{\mathcal{U}}$ while scanning this u-edge. Now, suppose that this is true for $i-1$. It follows from the inductive argument in Lemma 5.1 that the Vizing fans currently in $\mathcal{F}$ are all $\operatorname*{\mathcal{U}}$-avoiding. If we do not change $\chi$ or $\operatorname*{\mathcal{U}}$ while scanning $\operatorname*{\boldsymbol{e}}_{i}$, then we are done. Thus, we assume that this is not the case. The subroutine can modify $\chi$ and $\operatorname*{\mathcal{U}}$ in the following ways:
  • Rotating colors around some $\operatorname*{\mathbf{F}}_{j}$ so that $(u_{j},u_{i})$ is uncolored, setting $\chi(u_{j},u_{i})\leftarrow\alpha$ and removing $\operatorname*{\boldsymbol{e}}_{i}$ and $\operatorname*{\boldsymbol{e}}_{j}$ from $\operatorname*{\mathcal{U}}$.
  • Rotating colors around $\operatorname*{\mathbf{F}}_{i}$ so that $(u_{i},u_{j})$ is uncolored for some $\operatorname*{\mathbf{F}}_{j}\in\mathcal{F}$, setting $\chi(u_{j},u_{i})\leftarrow\alpha$ and removing $\operatorname*{\boldsymbol{e}}_{i}$ and $\operatorname*{\boldsymbol{e}}_{j}$ from $\operatorname*{\mathcal{U}}$.
  • Rotating colors around $\operatorname*{\mathbf{F}}_{i}$ and some $\operatorname*{\mathbf{F}}_{j}\in\mathcal{F}$ so that $(u_{i},u_{j})$ is uncolored, removing $\operatorname*{\boldsymbol{e}}_{i}$ and $\operatorname*{\boldsymbol{e}}_{j}$ from $\operatorname*{\mathcal{U}}$ and adding a u-fan $\operatorname*{\boldsymbol{f}}$ to $\operatorname*{\mathcal{U}}$.
In the first case, it follows from Lemma 3.9 (since the Vizing fans are all $\operatorname*{\mathcal{U}}$-avoiding) that rotating colors around $\operatorname*{\mathbf{F}}_{j}$ does not damage any u-component in $\operatorname*{\mathcal{U}}\setminus\{\operatorname*{\boldsymbol{e}}_{j}\}$. Furthermore, since $\operatorname*{\mathcal{U}}$ is separable, setting $\chi(u_{j},u_{i})\leftarrow\alpha$ does not damage any u-component in $\operatorname*{\mathcal{U}}\setminus\{\operatorname*{\boldsymbol{e}}_{i},\operatorname*{\boldsymbol{e}}_{j}\}$. Since we remove $\operatorname*{\boldsymbol{e}}_{i}$ and $\operatorname*{\boldsymbol{e}}_{j}$ from $\operatorname*{\mathcal{U}}$, we ensure that $\operatorname*{\mathcal{U}}$ remains separable. The same argument extends to the second case.
Finally, for the third case, it follows from Lemma 3.9 that rotating colors around $\operatorname*{\mathbf{F}}_{i}$ and $\operatorname*{\mathbf{F}}_{j}$ does not damage an u-component in $\operatorname*{\mathcal{U}}\setminus\{\operatorname*{\boldsymbol{e}}_{i},\operatorname*{\boldsymbol{e}}_{j}\}$, so after removing $\operatorname*{\boldsymbol{e}}_{i}$ and $\operatorname*{\boldsymbol{e}}_{j}$ from $\operatorname*{\mathcal{U}}$ we have that $\operatorname*{\mathcal{U}}$ is separable. Since we have that $\alpha\in C_{\operatorname*{\mathcal{U}}}(u_{i})\cap C_{\operatorname*{\mathcal{U}}}(u_{j})$ after removing $\operatorname*{\boldsymbol{e}}_{i}$ and $\operatorname*{\boldsymbol{e}}_{j}$ from $\operatorname*{\mathcal{U}}$, we can see that $\operatorname*{\boldsymbol{f}}$ is indeed a u-fan and that $\operatorname*{\mathcal{U}}\cup\{\operatorname*{\boldsymbol{f}}\}$ is separable. [^9] Note that, before removing $\operatorname*{\boldsymbol{e}}_{i}$ and $\operatorname*{\boldsymbol{e}}_{j}$ from $\operatorname*{\mathcal{U}}$, $c_{\operatorname*{\boldsymbol{e}}_{i}}(u_{i})=c_{\operatorname*{\boldsymbol{e}}_{j}}(u_{j})=\alpha$.
Lemma 5.3.
Each time Prune-Vizing-Fans modifies the coloring $\chi$, it removes at most $2$ u-edges from $\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$ and either (1) extends the coloring $\chi$ to one more edge, or (2) adds a u-fan to $\operatorname*{\mathcal{U}}$.
Proof.
This follows immediately from the description of the subroutine. ∎
The following lemma shows that this subroutine can be implemented efficiently.
Lemma 5.4.
We can implement Prune-Vizing-Fans to run in $O(\Delta\lambda_{\alpha})$ time.
Proof.
By Lemma 3.8, we can construct each of the $\operatorname*{\mathcal{U}}$-avoiding Vizing fans $\operatorname*{\mathbf{F}}_{i}$ in $O(\Delta)$ time. As we scan through the u-edges $\operatorname*{\boldsymbol{e}}_{1},\dots,\operatorname*{\boldsymbol{e}}_{\ell}$, we can maintain the set of vertices $S$ that are contained in the Vizing fans in $\mathcal{F}$, along with pointers to the corresponding Vizing fan. This allows us to either verify that $\operatorname*{\mathbf{F}}_{i}$ is vertex-disjoint from the Vizing fans in $\mathcal{F}$ or to find the first vertex in $\operatorname*{\mathbf{F}}_{i}$ that is contained in one of these fans in $O(\Delta)$ time. We can then appropriately update the coloring $\chi$ and the collection $\operatorname*{\mathcal{U}}$ in $O(\Delta)$ time using the data structures outlined in Section 3.2. It follows that the entire subroutine takes $O(\Delta\lambda_{\alpha})$ time. ∎
5.2 The Subroutine Reduce-U-Edges
As input, the subroutine Reduce-U-Edges is given a graph $G$, a partial $(\Delta+1)$-edge coloring $\chi$ of $G$, a color $\alpha$, a separable collection $\operatorname*{\mathcal{U}}$ with $\lambda_{\alpha}$ $\alpha$-primed u-edges and a set $\mathcal{F}$ of vertex-disjoint $\operatorname*{\mathcal{U}}$-avoiding Vizing fans for all of the u-edges in $\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$. Let $m_{\alpha}$ denote the number of edges with color $\alpha$ in the input coloring $\chi$. For each u-edge $\operatorname*{\boldsymbol{e}}\in\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$, let $\operatorname*{\mathbf{F}}_{\operatorname*{\boldsymbol{e}}}$ denote its Vizing fan in $\mathcal{F}$. Similarly, for each $\operatorname*{\mathbf{F}}\in\mathcal{F}$, let $\operatorname*{\boldsymbol{e}}_{\operatorname*{\mathbf{F}}}$ denote its corresponding u-edge in $\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$. The subroutine then either extends the coloring to the u-edges in $\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$ or uses them to construct u-fans.
[…]
More formally, the subroutine begins by scanning through all of the trivial Vizing fans $\operatorname*{\mathbf{F}}\in\mathcal{F}$ and calling $\textnormal{{Vizing}}(\operatorname*{\mathbf{F}})$, removing the $\operatorname*{\mathbf{F}}$ from $\mathcal{F}$ and $\operatorname*{\boldsymbol{e}}_{\operatorname*{\mathbf{F}}}$ from $\operatorname*{\mathcal{U}}$, and removing any other u-component damaged by calling $\textnormal{{Vizing}}(\operatorname*{\mathbf{F}})$ from $\operatorname*{\mathcal{U}}$ (see Lemma 3.9). Let $L\leftarrow 0$, $S\leftarrow\varnothing$ and $P_{\operatorname*{\mathbf{F}}}\leftarrow\varnothing$ for all $\operatorname*{\mathbf{F}}\in\mathcal{F}$. The subroutine then proceeds in rounds which consist of updating each of the paths $P_{\operatorname*{\mathbf{F}}}$ while maintaining the following invariant.
Invariant 5.5.
After updating a path $P_{\operatorname*{\mathbf{F}}}$, the following hold:
  1. 1.
[…]
  2. 2.
    The prefix paths $\{P_{\operatorname*{\mathbf{F}}}\}_{\operatorname*{\mathbf{F}}\in\mathcal{F}}$ are all edge-disjoint.
  3. 3.
    The set $S$ is the union of all the edges in the prefix paths $\{P_{\operatorname*{\mathbf{F}}}\}_{\operatorname*{\mathbf{F}}\in\mathcal{F}}$.
  4. 4.
    The collection $\operatorname*{\mathcal{U}}$ is separable and $|\mathcal{F}|=|\mathcal{E}_{\alpha}(\operatorname*{\mathcal{U}})|$.
  5. 5.
    The Vizing fans in $\mathcal{F}$ are $\operatorname*{\mathcal{U}}$-avoiding and vertex-disjoint.
The prefix paths $\{P_{\operatorname*{\mathbf{F}}}\}_{\operatorname*{\mathbf{F}}\in\mathcal{F}}$ maintained by the algorithm also satisfy the following invariant.
Invariant 5.6.
At the start of each round, we have that $|P_{\operatorname*{\mathbf{F}}}|=L$ for all $\operatorname*{\mathbf{F}}\in\mathcal{F}$.
[…]
1 $P_{\operatorname*{\mathbf{F}}}\leftarrow\textnormal{{Vizing-Path}}(\operatorname*{\mathbf{F}})_{\leq|P_{\operatorname*{\mathbf{F}}}|+1}$
2 Let $(x,y)$ be the $(L+1)^{th}$ edge in $P_{\operatorname*{\mathbf{F}}}$ and $y$ be an endpoint of $P_{\operatorname*{\mathbf{F}}}$
3 if $(x,y)\in S$ then
4       Let $\operatorname*{\mathbf{F}}^{\prime}\in\mathcal{F}$ be the Vizing fan such that $(x,y)\in P_{\operatorname*{\mathbf{F}}^{\prime}}$
5       if $(x,y)$ appears in the same orientation in $P_{\operatorname*{\mathbf{F}}}$ and $P_{\operatorname*{\mathbf{F}}^{\prime}}$ then
6             Let $(z,x)$ and $(z^{\prime},x)$ be the edges appearing before $(x,y)$ in $P_{\operatorname*{\mathbf{F}}}$ and $P_{\operatorname*{\mathbf{F}}^{\prime}}$ respectively
7             $\beta\leftarrow\chi(z,x)$
8             $\chi(z,x)\leftarrow\bot$ and $\chi(z^{\prime},x)\leftarrow\bot$
9             $\textnormal{{Vizing}}(\operatorname*{\mathbf{F}})$
10             $\textnormal{{Vizing}}(\operatorname*{\mathbf{F}}^{\prime})$
11             Remove $\operatorname*{\mathbf{F}}$ and $\operatorname*{\mathbf{F}}^{\prime}$ from $\mathcal{F}$ and $\operatorname*{\boldsymbol{e}}_{\operatorname*{\mathbf{F}}}$ and $\operatorname*{\boldsymbol{e}}_{\operatorname*{\mathbf{F}}^{\prime}}$ from $\operatorname*{\mathcal{U}}$
12             Add the u-fan $\operatorname*{\boldsymbol{f}}=(x,z,z^{\prime},\beta,\alpha)$ to $\operatorname*{\mathcal{U}}$
13
14      else
15             $\chi(x,y)\leftarrow\bot$
16             $\textnormal{{Vizing}}(\operatorname*{\mathbf{F}})$
17             $\textnormal{{Vizing}}(\operatorname*{\mathbf{F}}^{\prime})$
18             Remove $\operatorname*{\mathbf{F}}$ and $\operatorname*{\mathbf{F}}^{\prime}$ from $\mathcal{F}$ and $\operatorname*{\boldsymbol{e}}_{\operatorname*{\mathbf{F}}}$ and $\operatorname*{\boldsymbol{e}}_{\operatorname*{\mathbf{F}}^{\prime}}$ from $\operatorname*{\mathcal{U}}$
19
20      $S\leftarrow S\setminus(P_{\operatorname*{\mathbf{F}}}\cup P_{\operatorname*{\mathbf{F}}^{\prime}})$
21       return
[…]
23 if $P_{\operatorname*{\mathbf{F}}}$ is maximal  then
24       $\textnormal{{Vizing}}(\operatorname*{\mathbf{F}})$
25       Remove $\operatorname*{\mathbf{F}}$ from $\mathcal{F}$ and $\operatorname*{\boldsymbol{e}}_{\operatorname*{\mathbf{F}}}$ from $\operatorname*{\mathcal{U}}$
26       $S\leftarrow S\setminus P_{\operatorname*{\mathbf{F}}}$
27       if there exists $\operatorname*{\boldsymbol{g}}\in\operatorname*{\mathcal{U}}$ with $c_{\operatorname*{\boldsymbol{g}}}(y)=\chi(x,y)$ then
28             Remove $\operatorname*{\boldsymbol{g}}$ from $\operatorname*{\mathcal{U}}$ (see Lemma 3.9)
29
30      if there exists $\operatorname*{\mathbf{F}}^{\prime}\in\mathcal{F}$ such that $y\in\operatorname*{\mathbf{F}}$ then
31             Remove $\operatorname*{\mathbf{F}}^{\prime}$ from $\mathcal{F}$ and $\operatorname*{\boldsymbol{e}}_{\operatorname*{\mathbf{F}}^{\prime}}$ from $\operatorname*{\mathcal{U}}$ // $\operatorname*{\boldsymbol{e}}_{F^{\prime}}$ might also get removed in Algorithm 4
32             $S\leftarrow S\setminus P_{\operatorname*{\mathbf{F}}^{\prime}}$
33
[…]
Analysis of Reduce-U-Edges
The following lemmas summarize the main properties of the subroutine Reduce-U-Edges.
Lemma 5.7.
The subroutine Reduce-U-Edges satisfies 5.5.
Proof.
[…]
After activating and removing all of the trivial Vizing fans in $\mathcal{F}$, Conditions 1-3 are clearly still satisfied (since the paths are all empty). For Condition 4, note that the subroutine removes any damaged u-components from $\operatorname*{\mathcal{U}}$, removes $\operatorname*{\boldsymbol{e}}$ from $\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$ if it removes $\operatorname*{\mathbf{F}}_{\operatorname*{\boldsymbol{e}}}$ from $\mathcal{F}$, and that activating some trivial $\operatorname*{\mathbf{F}}_{\operatorname*{\boldsymbol{e}}}\in\mathcal{F}$ cannot damage any $\operatorname*{\boldsymbol{e}}^{\prime}\in\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})\setminus\{\operatorname*{\boldsymbol{e}}\}$ since their Vizing fans are vertex-disjoint. For Condition 5, note that the subroutine does not add any u-component to $\operatorname*{\mathcal{U}}$, so the vertex-disjoint Vizing fans remain $\operatorname*{\mathcal{U}}$-avoiding. This establishes the base case of the induction.
[…]
Conditions 2 and 3: After extending $P_{\operatorname*{\mathbf{F}}}$ by one more edge, we check if there is some other path $P_{\operatorname*{\mathbf{F}}^{\prime}}$ that intersects the updated path $P_{\operatorname*{\mathbf{F}}}$ at this new edge. If so, we remove both $\operatorname*{\mathbf{F}}$ and $\operatorname*{\mathbf{F}}^{\prime}$ from $\mathcal{F}$, ensuring that the remaining paths are edge-disjoint. Otherwise, the paths are all edge-disjoint. It’s straightforward to verify that $S$ is updated correctly in each case.
Condition 4: We first show that the collection $\operatorname*{\mathcal{U}}$ remains separable. If this call to Update-Path does not change $\chi$ or $\operatorname*{\mathcal{U}}$, then clearly $\operatorname*{\mathcal{U}}$ remains separable. Note that the Vizing fans in $\mathcal{F}$ are all $\operatorname*{\mathcal{U}}$-avoiding and recall Lemma 3.9. We now consider the following three cases.
(1) If we enter the if statement on Algorithm 4, then calling $\textnormal{{Vizing}}(\operatorname*{\mathbf{F}})$ and $\textnormal{{Vizing}}(\operatorname*{\mathbf{F}}^{\prime})$ after uncoloring edges on their Vizing chains does not damage any u-component in $\operatorname*{\mathcal{U}}$ apart from $\operatorname*{\boldsymbol{e}}_{\operatorname*{\mathbf{F}}}$ and $\operatorname*{\boldsymbol{e}}_{\operatorname*{\mathbf{F}}^{\prime}}$, which are both removed from $\operatorname*{\mathcal{U}}$. This is because shifting colors around the Vizing fans $\operatorname*{\mathbf{F}}$ and $\operatorname*{\mathbf{F}}^{\prime}$ does not damage any u-components apart from $\operatorname*{\boldsymbol{e}}_{\operatorname*{\mathbf{F}}}$ and $\operatorname*{\boldsymbol{e}}_{\operatorname*{\mathbf{F}}^{\prime}}$ (see Lemma 3.9) and truncating an alternating path before flipping it ensures that we do not remove any colors from the palettes of the vertices on that path. Finally, note that the u-fan $\operatorname*{\boldsymbol{f}}$ that we add to $\operatorname*{\mathcal{U}}$ only uses colors that were previously unavailable. (2) If we enter the else statement on Algorithm 4, then the argument is completely analogous to the previous case, except that we do not add any u-components to $\operatorname*{\mathcal{U}}$ (see Lemma 3.9). (3) If we enter the if statement on Algorithm 4, then calling $\textnormal{{Vizing}}(\operatorname*{\mathbf{F}})$ damages at most one u-component in $\operatorname*{\mathcal{U}}$ apart from $\operatorname*{\boldsymbol{e}}_{\operatorname*{\mathbf{F}}}$ (see Lemma 3.9). In particular, if there is such a u-component $\operatorname*{\boldsymbol{g}}$, it will contain the vertex $y$ and have $c_{\operatorname*{\boldsymbol{g}}}(x)=\chi(x,y)$. Thus, we check if such a u-component exists and remove it from $\operatorname*{\mathcal{U}}$.
Finally, to see that $|\mathcal{F}|=|\mathcal{E}_{\alpha}(\operatorname*{\mathcal{U}})|$, we can verify that we remove $\operatorname*{\boldsymbol{e}}$ from $\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$ if and only if we remove $\operatorname*{\mathbf{F}}_{\operatorname*{\boldsymbol{e}}}$ from $\mathcal{F}$. It’s clear that whenever we remove some $\operatorname*{\mathbf{F}}^{\prime}$ from $\mathcal{F}$, we also remove $\operatorname*{\boldsymbol{e}}_{\operatorname*{\mathbf{F}}^{\prime}}$ from $\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$. Similarly, if we remove a u-edge $\operatorname*{\boldsymbol{e}}$ from $\mathcal{E}_{\alpha}(\operatorname*{\mathcal{U}})$ in Algorithm 4, then we can see that we then remove $\operatorname*{\mathbf{F}}_{\operatorname*{\boldsymbol{e}}}$ from $\mathcal{F}$ in Algorithm 4 immediately afterwards.
Condition 5: Let $\operatorname*{\mathbf{F}}^{\prime}$ be some Vizing fan that remains in $\mathcal{F}$ after handling this call. Then, for any Vizing fan $\operatorname*{\mathbf{F}}^{\prime\prime}$ that was activated during the call, we know that its corresponding Vizing chain did not end at $\operatorname*{\mathbf{F}}^{\prime}$, otherwise $\operatorname*{\mathbf{F}}^{\prime}$ would have been removed from $\mathcal{F}$ (see Algorithm 4). Thus, activating $\operatorname*{\mathbf{F}}^{\prime\prime}$ (which is vertex-disjoint from $\operatorname*{\mathbf{F}}^{\prime}$) does not change the palettes of any vertices in $\operatorname*{\mathbf{F}}^{\prime}$. Note that, since the center of $\operatorname*{\mathbf{F}}^{\prime}$ is missing $\alpha$, it is not possible for the Vizing chain of $\operatorname*{\mathbf{F}}^{\prime\prime}$ to contain any edges in $\operatorname*{\mathbf{F}}^{\prime}$ without ending at $\operatorname*{\mathbf{F}}^{\prime}$. Hence, the Vizing fans remaining in $\mathcal{F}$ are still valid Vizing fans and remain vertex-disjoint. To see that they also remain $\operatorname*{\mathcal{U}}$-avoiding, note that if we add a u-fan $\operatorname*{\boldsymbol{f}}$ to $\operatorname*{\mathcal{U}}$ (see Algorithm 4), the color that $\operatorname*{\boldsymbol{f}}$ assigns to a vertex $x\in\operatorname*{\boldsymbol{f}}$ was previously not available at $x$ before the start of the call to Update-Path. Thus, if $x$ is contained in the Vizing fan $\operatorname*{\mathbf{F}}^{\prime}$, its color within $\operatorname*{\mathbf{F}}^{\prime}$ is not $c_{\operatorname*{\boldsymbol{f}}}(x)$.
[…]
Lemma 5.8.
The subroutine Reduce-U-Edges satisfies 5.6.
Proof.
[…]
Lemma 5.9.
Each time Reduce-U-Edges modifies the coloring $\chi$, it removes at most $2$ u-edges from $\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$ and either (1) extends the coloring $\chi$ to one more edge and removes at most one other u-component from $\operatorname*{\mathcal{U}}$, or (2) adds a u-fan to $\operatorname*{\mathcal{U}}$.
Proof.
[…]
Running Time: We now show how to implement this subroutine and analyse its running time. We begin with the following claim which shows that Update-Path can be implemented efficiently.
Claim 5.10.
[…]
Proof.
[…]
If this call does not remove $\operatorname*{\mathbf{F}}$ from $\mathcal{F}$, then we know that $P_{\operatorname*{\mathbf{F}}}$ is not maximal and is also edge-disjoint from the other prefix paths maintained by the subroutine. In this case, we update $S$ in $O(1)$ time and do not modify $\chi$ or $\operatorname*{\mathcal{U}}$.
On the other hand, if this call does remove $\operatorname*{\mathbf{F}}$ from $\mathcal{F}$, then we need to activate $O(1)$ many Vizing chains of length $O(L)$, which can be done in $O(\Delta+L)$ time. Removing the edges of the corresponding paths from $S$ can also be done in $O(L)$ time. Finally, using the data structures outlined in Section 3.2, we can update $\mathcal{F}$ and $\operatorname*{\mathcal{U}}$ in $O(1)$ time. ∎
Let $\mathcal{F}^{\star}$ denote the set $\mathcal{F}$ when we first call the subroutine Reduce-U-Edges. For each $\operatorname*{\mathbf{F}}\in\mathcal{F}^{\star}$, let $L_{\operatorname*{\mathbf{F}}}$ denote the value of $L$ at the time that $\operatorname*{\mathbf{F}}$ is removed from $\mathcal{F}$.
Claim 5.11.
The total time spent handling calls to Update-Path is at most
$O(\Delta\lambda_{\alpha})+O(1)\cdot\sum_{\operatorname*{\mathbf{F}}\in\mathcal{F}^{\star}}L_{\operatorname*{\mathbf{F}}}.$ (1)
Proof.
[…]
$\sum_{\operatorname*{\mathbf{F}}\in\mathcal{F}^{\star}}O(\Delta+L_{\operatorname*{\mathbf{F}}})\leq O(\Delta\lambda_{\alpha})+O(1)\cdot\sum_{\operatorname*{\mathbf{F}}\in\mathcal{F}^{\star}}L_{\operatorname*{\mathbf{F}}}.\qed$
Now, assume w.l.o.g. that =$\mathcal{F}^{\star}=\{\operatorname*{\mathbf{F}}_{1},\dots,\operatorname*{\mathbf{F}}_{\lambda_{\alpha}}\}$ and $L_{\lambda_{\alpha}}\leq\dots\leq L_{1}$, where $L_{i}:=L_{\operatorname*{\mathbf{F}}_{i}}$. Recall that $m_{\alpha}$ denotes the number of edges with color $\alpha$ when we first call the subroutine.
Claim 5.12.
For each $i\in[\lambda_{\alpha}]$, we have that $L_{i}\leq 4(m_{\alpha}+2\lambda_{\alpha})/i$.
Proof.
Let $i\in[\lambda_{\alpha}]$, $\tau=4(m_{\alpha}+2\lambda_{\alpha})/i$ and suppose that $\tau<L_{i}$. Then, at the start of iteration $\tau+1$, the Vizing fans $\operatorname*{\mathbf{F}}_{1},\dots\operatorname*{\mathbf{F}}_{i}$ are all contained in $\mathcal{F}$. Furthermore, it follows from Invariants 5.5 and 5.6 that, at the start of iteration $\tau+1$, the alternating paths $P_{\operatorname*{\mathbf{F}}_{1}},\dots,P_{\operatorname*{\mathbf{F}}_{i}}$ form a collection of $i$ edge-disjoint $\{\alpha,\cdot\}$-alternating paths of length $\tau$ in $G$. Let $T$ denote the total length of these paths. We can observe that $T$ is at most $3m_{\alpha}^{\prime}$, where $m_{\alpha}^{\prime}$ is the number of edges that currently have color $\alpha$, since at least a third of the edges in each of these paths has color $\alpha$ (note that any $\{\alpha,\cdot\}$-alternating path of length $k$ has at least $\lfloor k/2\rfloor$ edges with color $\alpha$). We can also observe that $m_{\alpha}^{\prime}\leq m_{\alpha}+2\lambda_{\alpha}$ since the number of edges with color $\alpha$ increases by at most $2$ each time we activate a Vizing fan in $\mathcal{F}$. Thus, it follows that
$4(m_{\alpha}+2\lambda_{\alpha})=i\tau\leq T\leq 3m_{\alpha}^{\prime}\leq 3(m_{\alpha}+2\lambda_{\alpha}),$
giving a contradiction. ∎
Lemma 5.13.
We can implement Reduce-U-Edges to run in $O((m_{\alpha}+\Delta\lambda_{\alpha})\log n)$ time.
Proof.
Activating the trivial Vizing fans in $\mathcal{F}$ when the subroutine is first called can be done in $O(\Delta\lambda_{\alpha})$ time. The running time of the rest of the subroutine is dominated by the time taken to handle the call to Update-Path. Combining Claims 5.11 and 5.12, it follows that the total time spent handling calls to Update-Path is at most
$O(\Delta\lambda_{\alpha})+O(1)\cdot\sum_{\operatorname*{\mathbf{F}}\in\mathcal{F}^{\star}}L_{\operatorname*{\mathbf{F}}}\leq O(\Delta\lambda_{\alpha})+O(1)\cdot\sum_{i=1}^{\lambda_{\alpha}}\frac{m_{\alpha}+\lambda_{\alpha}}{i}$
$\leq O((m_{\alpha}+\Delta\lambda_{\alpha})\log n).\qed$
5.3 Analysis of Construct-U-Fans (Proof of Lemma 4.1)
The algorithm begins by constructing a separable collection of $\lambda$ u-components $\operatorname*{\mathcal{U}}$. The following lemma shows that this can be done efficiently.
Lemma 5.14.
Given a graph $G$ and a partial $(\Delta+1)$-edge coloring $\chi$ of $G$ with $\lambda$ uncolored edges, we can construct a separable collection of $\lambda$ u-edges $\operatorname*{\mathcal{U}}$ in $O(m)$ time.
Proof.
The algorithm first initializes an empty separable collection $\operatorname*{\mathcal{U}}=\varnothing$. It then scans through the edges of the graph and retrieves the $\lambda$ edges $e_{1},\dots,e_{\lambda}$ that are uncolored by $\chi$. The algorithm then scans through each of these uncolored edges and, for each $e_{i}=(u_{i},v_{i})$, picks a missing color $\alpha_{i}\in\mathsf{miss}_{\chi}(u_{i})\setminus C_{\operatorname*{\mathcal{U}}}(u_{i})$ (see 3.5) and adds the u-edge $\operatorname*{\boldsymbol{e}}_{i}:=(u_{i},v_{i},\alpha_{i})$ to $\operatorname*{\mathcal{U}}$. It’s easy to verify that the resulting collection $\operatorname*{\mathcal{U}}$ is separable and contains $\lambda$ u-edges. Using the data structures outlined in Section 3.2, finding such a color and adding a u-edge to $\operatorname*{\mathcal{U}}$ can be done in $O(1)$ time. Thus, this entire process can be implemented in $O(m)$ time. ∎
Let $m_{\alpha}$ denote the number of edges that have color $\alpha$ w.r.t. the initial coloring $\chi$ when we first call Construct-U-Fans. The following claim shows that we cannot create too many more edges with the color $\alpha$ throughout this sequence of calls to the subroutines Prune-Vizing-Fans and Reduce-U-Edges for each color $\alpha\in[\Delta+1]$.
Claim 5.15.
For each color $\alpha\in[\Delta+1]$, we have that at most $m_{\alpha}+O(\lambda)$ edges have color $\alpha$ throughout the entire run of the algorithm Construct-U-Fans.
Proof.
This follows from the fact that the number of edges with color $\alpha$ can only increase when we extend the coloring $\chi$ to some uncolored edge. Furthermore, it can only increase by at most $2$ every time this happens. Thus, for any $\alpha\in[\Delta+1]$, the number of edges with color $\alpha$ can increase by at most $\lambda$ throughout the entire run of the algorithm Construct-U-Fans. ∎
Let $\lambda_{\alpha}$ denote the number of $\alpha$-primed u-edges in the initial collection $\operatorname*{\mathcal{U}}$. Note that the number of such u-edges can only decrease throughout the run of this algorithm. It follows from Lemmas 5.4 and 5.13 that the total running time of the algorithm Construct-U-Fans across all of these calls to the subroutines Prune-Vizing-Fans and Reduce-U-Edges is at most
$\sum_{\alpha\in[\Delta+1]}\left(O(\lambda_{\alpha}\Delta)+O((m_{\alpha}+\lambda+\lambda_{\alpha}\Delta)\log n)\right)\leq O((m+\Delta\lambda)\log n).$
Lemma 5.16.
The algorithm Construct-U-Fans either extends the coloring $\chi$ to at least $\lambda/9$ more edges or returns a separable collection of at least $\lambda/9$ u-fans $\operatorname*{\mathcal{U}}$.
Proof.
To see why this lemma holds, consider the following three quantities and how they evolve over time throughout the execution of the algorithm Construct-U-Fans: The number of edges that the algorithm has extended the coloring to so far, $\Psi_{\texttt{c}}$, the number of u-edges in $\operatorname*{\mathcal{U}}$, $\Psi_{\texttt{e}}:=|\mathcal{E}|$, and the number of u-fans in $\operatorname*{\mathcal{U}}$, $\Psi_{\texttt{f}}:=|\mathcal{U}\setminus\mathcal{E}|$. Immediately after constructing the separable collection of $\lambda$ u-edges $\operatorname*{\mathcal{U}}$, it holds that $\Psi_{\texttt{c}}=0$, $\Psi_{\texttt{e}}=\lambda$ and $\Psi_{\texttt{f}}=0$. We next argue that by the time the execution of the algorithm has finished, either $\Psi_{\texttt{c}}\geq\lambda/9$ or $\Psi_{\texttt{f}}\geq\lambda/9$ must hold. To this end, we employ the following potential function argument.
Claim 5.17.
The quantity $3(\Psi_{\textnormal{{f}}}+2\Psi_{\textnormal{{c}}})+\Psi_{\textnormal{{e}}}$ is non-decreasing throughout the algorithm’s execution.
Proof.
Consider first a call made by algorithm Construct-U-Fans to the Prune-Vizing-Fans subroutine. By Lemma 5.3, the subroutine repeatedly removes at most $2$ u-edges from $\operatorname*{\mathcal{U}}$ and either (1) adds a u-fan to $\operatorname*{\mathcal{U}}$, or (2) extends the coloring to another edge. This may decrease $\Psi_{\texttt{e}}$ by at most $2$ while increasing one of $\Psi_{\texttt{f}}$ or $\Psi_{\texttt{c}}$ by at least $1$, so $3(\Psi_{\texttt{f}}+2\Psi_{\texttt{c}})+\Psi_{\texttt{e}}$ increases in this case.
Next, consider a call to the Reduce-U-Edges subroutine. By Lemma 5.9, during any iteration, the subroutine removes at most $3$ u-edges from $\operatorname*{\mathcal{U}}$ and either (1) adds a u-fan to $\operatorname*{\mathcal{U}}$, or (2) extends the coloring to another edge while removing at most one u-fan from $\operatorname*{\mathcal{U}}$. This may decrease $\Psi_{\texttt{e}}$ by at most $3$ while increasing $\Psi_{\texttt{f}}+2\Psi_{\texttt{c}}$ by at least $1$, so $3(\Psi_{\texttt{f}}+2\Psi_{\texttt{c}})+\Psi_{\texttt{e}}$ cannot decrease in value. ∎
Immediately after constructing the separable collection of $\lambda$ u-edges in $\operatorname*{\mathcal{U}}$, we have $3(\Psi_{\texttt{f}}+2\Psi_{\texttt{c}})+\Psi_{\texttt{e}}=\lambda$. 5.17 implies that $3(\Psi_{\texttt{f}}+2\Psi_{\texttt{c}})+\Psi_{\texttt{e}}\geq\lambda$ holds at all times afterwards. Since the algorithm iterates until $\mathcal{E}=\varnothing$, we know that $\Psi_{\texttt{e}}=0$ when the algorithm’s execution terminates, so at that moment we have $3(\Psi_{\texttt{f}}+2\Psi_{\texttt{c}})\geq\lambda$. Hence, either $\Psi_{\texttt{f}}\geq\lambda/9$ or $\Psi_{\texttt{c}}\geq\lambda/9$ must hold. ∎
6 The Algorithm Color-U-Fans (Proof of Lemma 4.2)
As input, the algorithm Color-U-Fans is given a graph $G$, a partial $(\Delta+1)$-edge coloring $\chi$ of $G$, and a collection of separable u-fans $\mathcal{U}$ of size $\lambda$. It then uses two further subroutines, Prime-U-Fans and Activate-U-Fans, to prime u-fans with the same colors and then activate them. More specifically, the algorithm first identifies the two least common colors $\alpha,\beta\in[\Delta+1]$. It calls the subroutine Prime-U-Fans which proceeds to prime $\Omega(\lambda/\Delta)$ of the u-fans in $\operatorname*{\mathcal{U}}$ with the colors $\alpha$ and $\beta$ (i.e. to modify the coloring $\chi$ and these u-fans so that they are $\{\alpha,\beta\}$-primed). It then calls Activate-U-Fans which extends the coloring $\chi$ to $\Omega(\lambda/\Delta)$ of the edges in these u-fans. The algorithm repeats this process until it has colored $\Omega(\lambda)$ edges. Algorithm 5 gives the pseudocode for Color-U-Fans.
1 for $\Delta/2$ iterations do
2       Let $\alpha,\beta\in[\Delta+1]$ be the two least common colors in $\chi$
[…]
5
[…]
Organization of Section 6: We begin by describing and analyzing the subroutines Prime-U-Fans and Activate-U-Fans used by Color-U-Fans before proving Lemma 4.2 in Section 6.3.
6.1 The Subroutine Prime-U-Fans
As input, this subroutine is given a graph $G$, a partial $(\Delta+1)$-edge coloring $\chi$ of $G$, a collection of separable u-fans $\mathcal{U}$ of size $\lambda$, and the two least common colors $\alpha,\beta\in[\Delta+1]$ (see 6.1). It then proceeds in iterations, where in each iteration it samples a u-fan $\operatorname*{\boldsymbol{f}}$ from $\mathcal{U}$ uniformly at random and attempts to prime $\operatorname*{\boldsymbol{f}}$ with the colors $\alpha$ and $\beta$. The subroutine maintains a subset $\Phi\subseteq\operatorname*{\mathcal{U}}$ of $\{\alpha,\beta\}$-primed u-fans. The subroutine performs iterations until $|\Phi|=\Omega(\lambda/\Delta)$, after which we proceed to call the subroutine Activate-U-Fans to extend the coloring $\chi$ to edges in the u-fans in $\Phi$. The pseudocode in Algorithm 6 gives a formal description of the subroutine Prime-U-Fans.
1 $\Phi\leftarrow\varnothing$ and $\lambda\leftarrow|\mathcal{U}|$
2 while $|\Phi|<\lambda/(48\Delta)$ do
3       Sample $\operatorname*{\boldsymbol{f}}=(u,v,w,\gamma,\delta)\sim U$ independently and u.a.r.
4       $(\alpha^{\prime},\beta^{\prime})\leftarrow(\alpha,\beta)$     // $\alpha^{\prime}$ and $\beta^{\prime}$ ensure we never flip an $\{\alpha,\beta\}$-alt. path
5       if $\gamma=\beta$ or $\delta=\alpha$ then
6             $(\alpha^{\prime},\beta^{\prime})\leftarrow(\beta,\alpha)$
7
8      Let $P_{u}$ be the $\{\alpha^{\prime},\gamma\}$-alternating path starting at $u$
9       Let $P_{v}$ and $P_{w}$ be the $\{\beta^{\prime},\delta\}$-alternating paths starting at $v$ and $w$ respectively
10       $\textnormal{{cost}}(\operatorname*{\boldsymbol{f}}):=|P_{u}|+|P_{v}|+|P_{w}|$
11       if $\textnormal{{cost}}(\operatorname*{\boldsymbol{f}})\leq 128m/\lambda$ then
12             Let $S$ denote the set of endpoints of $P_{u}$, $P_{v}$ and $P_{w}$
13             if the vertices in $S\cup\{u,v,w\}$ are not in any u-fan in $\Phi$ then
14                   Flip the alternating paths $P_{u}$, $P_{v}$ and $P_{w}$
15                   Remove $\operatorname*{\boldsymbol{f}}$ and any damaged u-fans from $\mathcal{U}$
16                   Add the u-fan $(u,v,w,\alpha^{\prime},\beta^{\prime})$ to $\operatorname*{\mathcal{U}}$ and $\Phi$
17
18return $(\alpha,\beta)$
[…]
Analysis of Prime-U-Fans
We say that an iteration of (the while loop in) Prime-U-Fans is successful if it adds a u-fan to the set $\Phi$. Otherwise, we say that the iteration fails. Note that the algorithm repeatedly performs iterations until it has performed at least $\lambda/(48\Delta)$ successful iterations. The following lemmas summarize the main properties of the subroutine Prime-U-Fans.
Claim 6.1.
The total number of edges with color $\alpha$ or $\beta$ is $O(m/\Delta)$ at any point throughout the run of Prime-U-Fans.
Proof.
Since $\alpha$ and $\beta$ are initially the two least common colors, there are initially $O(m/\Delta)$ such edges. Since each successful iteration increases the number of such edges by $O(1)$ and there are at most $O(\lambda/\Delta)\leq O(m/\Delta)$ such iterations, the claim follows. ∎
Claim 6.2.
$|\mathcal{U}|\geq(1-1/(2\Delta))\lambda$ at any point throughout the run of Prime-U-Fans.
Proof.
Consider a successful iteration of the algorithm where we sample a u-fan $\operatorname*{\boldsymbol{f}}\in\mathcal{U}$. The algorithm removes $\operatorname*{\boldsymbol{f}}$ from $\mathcal{U}$, along with any other damaged u-fans in $\mathcal{U}$. It follows from Lemma 3.4 that flipping the colors of the alternating paths $P_{u}$, $P_{v}$ and $P_{w}$ damages at most $6$ u-fans in $\operatorname*{\mathcal{U}}$. Since our algorithm runs for at most $\lambda/(48\Delta)$ iterations, it follows that $|\mathcal{U}|\geq\lambda-7\cdot\lambda/(48\Delta)$. ∎
Lemma 6.3.
The u-fans in $\Phi$ are all vertex-disjoint and $\{\alpha,\beta\}$-primed at any point throughout the run of Prime-U-Fans.
Proof.
If the algorithm samples a u-fan $\operatorname*{\boldsymbol{f}}$ that shares a vertex $x$ with a u-fan $\operatorname*{\boldsymbol{f}}^{\prime}\in\Phi$, then the iteration fails. Thus, the u-fans in $\Phi$ are all vertex-disjoint.
Now, note that whenever we add a u-fan $(u,v,w,\alpha^{\prime},\beta^{\prime})$ to $\Phi$, we do this immediately after flipping the paths $P_{u}$, $P_{v}$, and $P_{w}$. Thus, we have that $\alpha^{\prime}\in\mathsf{miss}_{\chi}(u)$ and $\beta^{\prime}\in\mathsf{miss}_{\chi}(v)\cap\mathsf{miss}_{\chi}(w)$, so the u-fan is $\{\alpha,\beta\}$-primed when we add it to $\Phi$. Since we only flip the paths in Algorithm 6 if their endpoints do not touch any u-fans in $\Phi$, this operation cannot change what colors are available at vertices in u-fans in $\Phi$, and hence cannot change whether or not any u-fan in $\Phi$ is $\{\alpha,\beta\}$-primed. ∎
The following standard claim bounds the total length of all maximal $\{c,\cdot\}$-alternating paths.
Claim 6.4.
For any color $c\in[\Delta+1]$, the total length of all maximal $\{c,\cdot\}$-alternating paths in $\chi$ is at most $4m$.
Proof.
Let $\mathcal{P}_{c}$ denote the set of all such alternating paths. First note that the total length of all alternating paths in $\mathcal{P}_{c}$ with length $1$ is at most $m$. Now, let $c^{\prime}\in[\Delta+1]\setminus\{c\}$ and let $P$ be a maximal $\{c,c^{\prime}\}$-alternating path with $|P|\geq 2$. We can observe that at least a third of the edges in $P$ have color $c^{\prime}$, and that each edge with color $c^{\prime}$ only appears in one path in $\mathcal{P}_{c}$. Thus, we have that
$\sum_{P\in\mathcal{P}_{c}}|P|\leq m+3\cdot|\{e\in E\mid\chi(e)\neq c\}|\leq 4m.\qed$
Lemma 6.5.
Each iteration of Prime-U-Fans succeeds with probability at least $1/4$.
Proof.
Consider the state of the algorithm at the start of some iteration. For each u-fan $\operatorname*{\boldsymbol{f}}^{\prime}\in\operatorname*{\mathcal{U}}$, let $P_{x}(\operatorname*{\boldsymbol{f}}^{\prime})$ denote the alternating path starting at $x\in\operatorname*{\boldsymbol{f}}^{\prime}$ that is considered by the algorithm if it samples $\operatorname*{\boldsymbol{f}}^{\prime}\in\operatorname*{\mathcal{U}}$. We define the cost of the u-fan $\operatorname*{\boldsymbol{f}}^{\prime}$ to be
$\textnormal{{cost}}(\operatorname*{\boldsymbol{f}}{\!{}^{\prime}}):=\sum_{x\in\operatorname*{\boldsymbol{f}}^{\prime}}|P_{x}(\operatorname*{\boldsymbol{f}}{\!{}^{\prime}})|.$
Let $\mathcal{U}^{\star}\subseteq\mathcal{U}$ denote the subset of u-fans $\operatorname*{\boldsymbol{f}}^{\prime}\in\operatorname*{\mathcal{U}}$ such that none of the alternating paths $\{P_{x}(\operatorname*{\boldsymbol{f}}^{\prime})\}_{x\in\operatorname*{\boldsymbol{f}}^{\prime}}$ have endpoints at some u-fan in $\Phi$. We can see that the iteration is successful if and only if the u-fan $\operatorname*{\boldsymbol{f}}\in\operatorname*{\mathcal{U}}$ sampled during the iteration satisfies $\textnormal{{cost}}(\operatorname*{\boldsymbol{f}})\leq 128m/\lambda$ (see Algorithm 6) and $\operatorname*{\boldsymbol{f}}\in\operatorname*{\mathcal{U}}^{\star}$ (see Algorithm 6). Thus, we now show that this happens with probability at least $1/4$.
We first begin with the following claim.
Claim 6.6.
Let $P$ be an $\{\alpha,\cdot\}$- or $\{\beta,\cdot\}$-alternating path in $\chi$. Then there are at most $2$ u-fans $\operatorname*{\boldsymbol{f}}^{\prime}\in\mathcal{U}$ that, if sampled during the iteration, might cause the algorithm to flip $P$.
Proof.
Suppose that the path $P$ is a $\{c,c^{\prime}\}$-alternating path for some colors $c\in\{\alpha,\beta\}$ and $c^{\prime}\notin\{\alpha,\beta\}$ with endpoints $x$ and $y$ [^10] Note that the algorithm never flips $\{\alpha,\beta\}$-alternating paths. Since the collection of u-fans $\mathcal{U}$ is separable, we know that at most one u-fan in $\mathcal{U}$ containing $x$ (resp. $y$) has $c^{\prime}$ assigned as the color missing at $x$ (resp. $y$). Thus, we must sample one of these $2$ u-fans for the algorithm to flip the path $P$. ∎
Let $\mathcal{P}$ denote the set of all maximal $\{\alpha,\cdot\}$- and $\{\beta,\cdot\}$-alternating paths in $\chi$. We can see that the expected cost of the u-fan $\operatorname*{\boldsymbol{f}}$ sampled during the iteration is
$\mathbb{E}[\textnormal{{cost}}(\operatorname*{\boldsymbol{f}})]=\frac{1}{|\mathcal{U}|}\cdot\sum_{\operatorname*{\boldsymbol{f}}\in\operatorname*{\mathcal{U}}}\sum_{x\in\operatorname*{\boldsymbol{f}}}|P_{x}(\operatorname*{\boldsymbol{f}})|\leq\frac{2}{|\mathcal{U}|}\cdot\sum_{P\in\mathcal{P}}|P|\leq\frac{32m}{\lambda},$
where we are using the facts that (1) each path in $P\in\mathcal{P}$ appears at most twice while summing over the paths $|P_{x}(\operatorname*{\boldsymbol{f}})|$ by 6.6, (2) that $|\mathcal{U}|\geq\lambda/2$ by 6.2, and (3) that that the total length of all paths in $\mathcal{P}$ is $8m$ by 6.4. Applying Markov’s inequality, it follows that
$\Pr\!\left[\textnormal{{cost}}(\operatorname*{\boldsymbol{f}})\geq\frac{128m}{\lambda}\right]\leq\frac{1}{4}.$
Since each u-fan in $\Phi$ has 3 vertices and at most $2\Delta$$\{\alpha,\cdot\}$- or $\{\beta,\cdot\}$-alternating paths end at each of these vertices, we know that there are at most $6\Delta|\Phi|$ alternating paths that could be considered during some iteration that end at a u-fan in $\Phi$. By 6.6, there are at most $12\Delta|\Phi|$ u-fans in $\mathcal{U}$ that could cause the algorithm to consider one of these paths. Thus, at least $|\mathcal{U}|-12\Delta|\Phi|$ of the u-fans in $\mathcal{U}$ are contained in $\mathcal{U}^{\star}$. It follows that
$|\mathcal{U}^{\star}|\geq|\mathcal{U}|-12\Delta|\Phi|\geq|\mathcal{U}|-\frac{\lambda}{4}\geq\frac{|\mathcal{U}|}{2},$
where we are using the facts that $|\Phi|\leq\lambda/(48\Delta)$ and $|\mathcal{U}|\geq\lambda/2$ from 6.2. Since $\operatorname*{\boldsymbol{f}}$ is sampled uniformly at random from $\operatorname*{\mathcal{U}}$, it follows that $\Pr[\operatorname*{\boldsymbol{f}}\in\mathcal{U}^{\star}]\geq 1/2$. The lemma follows by applying a union bound. ∎
Lemma 6.7.
Each iteration of Prime-U-Fans takes time $O(m/\lambda)$.
Proof.
Using standard data structures, each iteration can be implemented in time proportional to the length of the alternating paths $P_{u}$, $P_{v}$ and $P_{w}$ considered by the algorithm during the iteration. We can check if $|P_{u}|+|P_{v}|+|P_{w}|\leq 128m/\lambda$ in $O(m/\lambda)$ time by traversing these paths and aborting if we notice that their total length exceeds $128m/\lambda$. If their total length is at most $128m/\lambda$, then we can flip these paths and update the collection $\operatorname*{\mathcal{U}}$ in $O(m/\lambda)$ time. ∎
Lemma 6.8.
The subroutine Prime-U-Fans runs in time $O(m\log n/\Delta)$ with high probability.
Proof.
Since each iteration of Prime-U-Fans succeeds with probability at least $1/4$ by Lemma 6.5 and the subroutine performs iterations until it succeeds $\lambda/(48\Delta)$ times, it follows that it performs at most $O(\lambda\log n/\Delta)$ iterations with high probability. Since each iteration takes $O(m/\lambda)$ time by Lemma 6.7, it follows that the total running time is $O(m\log n/\Delta)$ with high probability. ∎
6.2 The Subroutine Activate-U-Fans
As input, this subroutine is given a graph $G$, a partial $(\Delta+1)$-edge coloring $\chi$ of $G$, and a subset $\Phi\subseteq\operatorname*{\mathcal{U}}$ of $\mu$ vertex-disjoint $\{\alpha,\beta\}$-primed u-fans such that at most $O(m/\Delta)$ edges have color $\alpha$ or $\beta$. The subroutine repeatedly picks any $\operatorname*{\boldsymbol{f}}\in\Phi$ and proceeds to activate the u-fan. It repeats this process until $\Phi=\varnothing$. The pseudocode in Algorithm 7 gives a formal description of the subroutine.
1 while $\Phi\neq\varnothing$ do
2       Let $\operatorname*{\boldsymbol{f}}\in\Phi$
3       Let $P$ and $P^{\prime}$ be the $\{\alpha,\beta\}$-alternating paths starting at the leaves of $\operatorname*{\boldsymbol{f}}$
4       Activate the u-fan $\operatorname*{\boldsymbol{f}}$ by flipping the path $P$ or $P^{\prime}$
5       Remove $\operatorname*{\boldsymbol{f}}$ and any damaged u-fans from $\operatorname*{\mathcal{U}}$ and $\Phi$
6
[…]
Analysis of Activate-U-Fans
The following lemmas summarise the key properties of the subroutine Activate-U-Fans.
Lemma 6.9.
The subroutine Activate-U-Fans extends the coloring $\chi$ to at least $\mu/2$ more edges.
Proof.
During each iteration of the while loop, we activate a u-fan $\operatorname*{\boldsymbol{f}}$ and extend the coloring to an uncolored edge in $\operatorname*{\boldsymbol{f}}$. By Lemma 3.4, this process damages at most $2$ u-fans in $\operatorname*{\mathcal{U}}$ (including $\operatorname*{\boldsymbol{f}}$). Thus, in each iteration $|\Phi|$ decreases by at most $2$, and hence we extend the coloring to at least $\mu/2$ more edges across all the iterations performed by the subroutine. ∎
Lemma 6.10.
The subroutine Activate-U-Fans has a running time of $O(m/\Delta)$.
Proof.
Let $\mathcal{P}$ denote the collection of maximal $\{\alpha,\beta\}$-alternating paths in the coloring $\chi$ when we first call Activate-U-Fans. Since at most $O(m/\Delta)$ edges have color $\alpha$ and $\beta$ in this coloring, we have that the total length of the paths in $\mathcal{P}$ is $O(m/\Delta)$ since the paths in $\mathcal{P}$ are all vertex-disjoint.
Throughout the run of Activate-U-Fans, we only modify $\chi$ by flipping the colors of $\{\alpha,\beta\}$-alternating paths (see Algorithm 7). Thus, the structure of the paths in $\mathcal{P}$ does not change (but their colors might be flipped). Furthermore, each time the subroutine flips an $\{\alpha,\beta\}$-alternating path, it removes any u-fan from $\Phi$ that contains an endpoint of this alternating path (see Algorithm 7). Thus, each path in $\mathcal{P}$ is only flipped at most once.
Using standard data structures, each iteration of the while loop can be implemented in time proportional to the length of the alternating path that is flipped during the iteration (see Algorithm 7). It follows that the total running time of the subroutine across all iterations is at most $\sum_{P\in\mathcal{P}}O(|P|)\leq O(m/\Delta)$. ∎
Claim 6.11.
$|\mathcal{U}|$ decreases by at most $2\mu$ throughout the run of Activate-U-Fans.
Proof.
During each iteration, we flip an alternating path and remove the damaged u-fans from $\operatorname*{\mathcal{U}}$ and $\Phi$. It follows from Lemma 3.4 that each iteration removes at most $2$ u-fans from these sets. Since we perform at most $\mu$ iterations, the claim follows. ∎
6.3 Analysis of Color-U-Fans (Proof of Lemma 4.2)
Given a separable collection of $\lambda$ u-fans $\mathcal{U}$, the algorithm Color-U-Fans repeatedly calls Prime-U-Fans and Activate-U-Fans with the set $\mathcal{U}$ as described in Algorithm 5. It repeats this process for $\Delta/2$ iterations. It follows from Claims 6.2 and 6.11 that $|\mathcal{U}|$ decreases by at most a $(1-1/\Delta)$ factor during each iteration. Thus, by Bernoulli’s inequality, we get that
$|\mathcal{U}|\geq\lambda\cdot\left(1-\frac{1}{\Delta}\right)^{\Delta/2}\geq\frac{\lambda}{2}$
throughout the entire run of the algorithm. In each iteration of Algorithm 5, we extend the coloring to $\Omega(\lambda/\Delta)$ edges in $O(m\log n/\Delta)$ time w.h.p. Thus, in total, we extend the coloring to $\Omega(\lambda)$ edges in $O(m\log n)$ time w.h.p.
7 Implementation and Data Structures
In this section, we describe the key data structures that we use to implement an edge coloring $\chi$ and a separable collection $\operatorname*{\mathcal{U}}$, allowing us to efficiently implement the operations performed by our algorithms. We first describe the data structures and then show how they can be used to efficiently implement the queries described in Section 3.2.
Implementing an Edge Coloring: Let $G=(V,E)$ and a graph of maximum degree $\Delta$ and let $\mathcal{C}:=[\Delta+1]\cup\{\bot\}$. We implement an edge coloring $\chi:E\longrightarrow\mathcal{C}$ of $G$ using the following:
  • The map $\phi:E\longrightarrow\mathcal{C}$ where $\phi(e):=\chi(e)$ for all $e\in E$.
  • The map $\phi^{\prime}:V\times\mathcal{C}\longrightarrow E$ where $\phi_{u}^{\prime}(c):=\{e\ni u\mid\chi(e)=c\}$.
  • The set $\chi^{-1}(c):=\{e\in E\mid\chi(e)=c\}$, for all $c\in\mathcal{C}$.
  • The set $\mathsf{miss}_{\chi}(u)\cap[\deg_{G}(u)+1]$, for all $u\in V$ [^11] We take this intersection with $[\deg_{G}(u)+1]$ instead of maintaining $\mathsf{miss}_{\chi}(u)$ directly to ensure that the space complexity and initialization time of the data structures are $\tilde{O}(m)$ and not $\Omega(\Delta n)$.
We implement all of the maps and sets using hashmaps, allowing us to perform membership queries, insertions and deletions in $O(1)$ time (see Proposition 3.6). [^12] By using balanced search trees instead of hashmaps, we can make the data structures deterministic while increasing the time taken to perform these operations to $O(\log n)$. The map $\phi^{\prime}$ allows us to check if a color $c\in[\Delta+1]$ is available at a vertex $u\in V$ in $O(1)$ time, and if it is not, to find the edge $e\ni u$ with $\chi(e)=c$. The sets $\{\chi^{-1}(c)\}_{c\in\mathcal{C}}$ allow us to easily return all edges with a specific color (including $\bot$). Furthermore, we can determine which color classes are the least common in $O(1)$ time. [^13] For example, we can then maintain a list of colors $c\in\mathcal{C}$ sorted by the values of $|\chi^{-1}(c)|$. Each time an edge $e$ changes color under $\chi$, we can easily update all of these data structures in $O(1)$ time. Furthermore, given $O(1)$ time query access to an edge coloring $\chi$, we can initialize these data structures in $O(m)$ time. We note that $\chi$ is a proper edge coloring if and only if $|\phi^{\prime}_{u}(c)|\leq 1$ for all $u\in V,c\in[\Delta+1]$.
Since the hashmap used to implement $\phi$ stores $m$ elements, it follows that it can be implemented with $O(m)$ space. Similarly, the map $\phi^{\prime}$ stores $2m$ elements (if $\{e\ni u\mid\chi(e)=c\}=\varnothing$, then we do not store anything for $\phi^{\prime}_{u}(c)$) and thus can be implemented with $O(m)$ space since each element has size $O(1)$ (recall that $|\phi^{\prime}_{u}(c)|\leq 1$ since the coloring is proper). since each set $\chi^{-1}(c)$ can be stored in space $O(|\chi^{-1}(c)|)$ and $\sum_{c}|\chi^{-1}(c)|=m$, these sets can be implemented in space $O(m)$. Similarly, since $\sum_{u}|\mathsf{miss}_{\chi}(u)\cap[\deg_{G}(u)+1]|=O(m)$, the sets $\mathsf{miss}_{\chi}(u)\cap[\deg_{G}(u)+1]$ can also be implemented in $O(m)$ space.
Implementing a Separable Collection: We implement a separable collection $\operatorname*{\mathcal{U}}$ in a similar manner using the following:
  • The map $\psi:V\times[\Delta+1]\longrightarrow\operatorname*{\mathcal{U}}$ where $\psi_{u}(c):=\{\operatorname*{\boldsymbol{g}}\in\operatorname*{\mathcal{U}}\mid u\in\operatorname*{\boldsymbol{g}},c_{\operatorname*{\boldsymbol{g}}}(u)=c\}$.
  • The set $C_{\operatorname*{\mathcal{U}}}(u):=\{c_{\operatorname*{\boldsymbol{g}}}(u)\mid\operatorname*{\boldsymbol{g}}\in\operatorname*{\mathcal{U}},u\in\operatorname*{\boldsymbol{g}}\}$, for all $u\in V$.
  • The set $\overline{C}_{\operatorname*{\mathcal{U}}}(u):=\left(\mathsf{miss}_{\chi}(u)\cap[\deg_{G}(u)+1]\right)\setminus C_{\operatorname*{\mathcal{U}}}(u)$, for all $u\in V$.
We again implement all of the maps and sets using hashmaps, allowing us to access and change entries in $O(1)$ time. We note that, since $\operatorname*{\mathcal{U}}$ is separable, $|\psi_{u}(c)|\leq 1$ for all $u\in V,c\in[\Delta+1]$. Thus, we can determine the size of $\operatorname*{\mathcal{U}}$ and also sample from $\operatorname*{\mathcal{U}}$ uniformly at random in $O(1)$ time. [^14] For example, we can sample a number $r\sim[|\mathcal{U}|]$ u.a.r. and then return the $r^{th}$ element in the hashmap that implements $\psi$. Each time we remove a color $c\in[\Delta+1]$ from the palette $\mathsf{miss}_{\chi}(u)$ of a vertex $u\in V$, we can update $\overline{C}_{\operatorname*{\mathcal{U}}}(u)$ in $O(1)$ time and check $\psi_{u}(c)$ in $O(1)$ time to find any u-component that has been damaged. Each time we add or remove a u-component from $\operatorname*{\mathcal{U}}$, we can update all of these data structures in $O(1)$ time. Furthermore, we can initialize these data structures for an empty collection in $O(m)$ time by creating an empty map $\psi$, empty sets $C_{\operatorname*{\mathcal{U}}}(u)$ for each $u\in V$ and copying the sets $\overline{C}_{\operatorname*{\mathcal{U}}}(u)=\mathsf{miss}_{\chi}(u)\cap[\deg_{G}(u)+1]$ for each $u\in V$ which are maintained by the data structures for the edge coloring $\chi$. Since $\operatorname*{\mathcal{U}}$ is separable, we can see that $\overline{C}_{\operatorname*{\mathcal{U}}}(u)\neq\varnothing$. Thus, whenever we want a color from the set $\mathsf{miss}_{\chi}(u)\setminus C_{\operatorname*{\mathcal{U}}}(u)$, it suffices to take an arbitrary color from $\overline{C}_{\operatorname*{\mathcal{U}}}(u)$.
For each $\operatorname*{\boldsymbol{g}}\in\operatorname*{\mathcal{U}}$, we can see that $\operatorname*{\boldsymbol{g}}$ is contained at most $3$ times in $\psi$. Thus, the total space required to store the hashmap that implements $\psi$ is $O(|\operatorname*{\mathcal{U}}|)$. Since $\operatorname*{\mathcal{U}}$ is separable, the u-components in $\operatorname*{\mathcal{U}}$ are edge-disjoint, and thus $|\mathcal{U}|\leq m$. It follows that map $\psi$ can be stored with $O(m)$ space. For each $u\in V$, we can observe that $|C_{\operatorname*{\mathcal{U}}}(u)|\leq\deg_{G}(u)$ since at most $\deg_{G}(u)$ many u-components in $\operatorname*{\mathcal{U}}$ contain the vertex $u$, and $|\overline{C}_{\operatorname*{\mathcal{U}}}(u)|\leq\deg_{G}(u)+1$ since $\overline{C}_{\operatorname*{\mathcal{U}}}(u)\subseteq[\deg_{G}(u)+1]$. Thus, the total space required to store the sets $\{C_{\operatorname*{\mathcal{U}}}(u)\}_{u\in V}$ and $\{\overline{C}_{\operatorname*{\mathcal{U}}}(u)\}_{u\in V}$ is $O(m)$.
7.1 Implementing the Operations from Section 3.2
We now describe how to implement each of the operations from Section 3.2.
Implementing $\textsc{Initialize}(G,\chi)$: Suppose that we are given the graph $G$ and $O(1)$ time query access to an edge coloring $\chi$ of $G$. We can initialize the data structures used to maintain the maps $\phi$ and $\phi^{\prime}$ in $O(m)$ time. We can then scan through the edges $e\in E$ and initialize the sets $\chi^{-1}(c)$ in $O(m)$ time. Finally, we can scan through the vertices $u\in V$ and initialize the sets $\mathsf{miss}_{\chi}(u)\cap[\deg_{G}(u)+1]$ in $O(m)$ time. We can then initialize the data structures for an empty separable collection in $O(m)$ time by creating an empty map $\psi$ and, for each $u\in V$, initializing the sets $C_{\operatorname*{\mathcal{U}}}(u)\leftarrow\varnothing$ and $\overline{C}_{\operatorname*{\mathcal{U}}}(u)\leftarrow\mathsf{miss}_{\chi}(u)\cap[\deg_{G}(u)+1]$.
Implementing $\textsc{Insert}_{\operatorname*{\mathcal{U}}}(\operatorname*{\boldsymbol{g}})$: By performing at most $3$ queries to the map $\psi$, we can check if $\operatorname*{\mathcal{U}}\cup\{\operatorname*{\boldsymbol{g}}\}$ is separable. If so we can update $\psi$ and the sets $C_{\operatorname*{\mathcal{U}}}(x)$ and $\overline{C}_{\operatorname*{\mathcal{U}}}(x)$ for $x\in\operatorname*{\boldsymbol{g}}$ in $O(1)$ time in order to insert $\operatorname*{\boldsymbol{g}}$ into $\operatorname*{\mathcal{U}}$. Otherwise, we return fail.
Implementing $\textsc{Delete}_{\operatorname*{\mathcal{U}}}(\operatorname*{\boldsymbol{g}})$: We can first make a query to $\psi$ to ensure that $\operatorname*{\boldsymbol{g}}\in\operatorname*{\mathcal{U}}$. If so, we can update $\psi$ and the sets $C_{\operatorname*{\mathcal{U}}}(x)$ and $\overline{C}_{\operatorname*{\mathcal{U}}}(x)$ for $x\in\operatorname*{\boldsymbol{g}}$ in $O(1)$ time to remove $\operatorname*{\boldsymbol{g}}$ from $\operatorname*{\mathcal{U}}$.
[…]
Acknowledgements
Part of this work was conducted while the first and second named authors were visiting the Simons Institute for the Theory of Computing as part of the Sublinear Algorithms program.
Appendix A Logarithmic Factors and Small Values of $\Delta$
The main contribution and the primary message of our work is presenting the first near-linear time algorithm for $(\Delta+1)$-edge coloring and polynomial time factor improvements over prior work. As such, in the presentation of our algorithm and its analysis, we made no attempt to optimize the logarithmic factors, which led to an algorithm with $O(m\log^{3}{n})$ time for finding a $(\Delta+1)$-edge coloring with high probability.
Nevertheless, we can indeed optimize our algorithm to achieve a runtime of only $O(m\log{\Delta})$ time (with high probability), which matches the classical bound of [CH82, COS01] for $\Delta$-coloring bipartite graphs (although that algorithm is deterministic). Our algorithm is also related to a recent line of work in [BD23, BD24, Dha24] that focuses on the $\Delta=n^{o(1)}$ case; our result improves prior $O(m\Delta^{4}\log\Delta)$ bounds of [BD24] exponentially in terms of dependence on $\Delta$.
Theorem A.1.
There is a randomized algorithm that, given any simple undirected graph $G$ on $n$ vertices and $m$ edges with maximum degree $\Delta$, finds a $(\Delta+1)$-edge coloring of $G$ in $O(m\log{\Delta})$ time with high probability.
Let us first start by listing where the $\log(n)$-terms come from in the proof of Theorem 1.1:
  1. 1.
    In Lemma 4.1, we are losing an $O(\log{n})$ factor in the subroutine Reduce-U-Edges while turning u-edges into u-fans or coloring them. Specifically, in the proof of Lemma 5.13, we lose an additional $O(\log{n})$ factor when handling all u-edges (see the Harmonic series in the proof).
  2. 2.
    In Lemma 4.2, when coloring u-fans, we are losing an $O(\log{n})$ factor to ensure the probabilistic guarantees of the algorithm hold with high probability in each step (see Lemma 6.8).
  3. 3.
    Our application of Lemmas 4.1 and 4.2 can color a constant fraction of remaining uncolored edges, hence, we need to run them $O(\log{n})$ times to color all uncolored edges (see Lemma 4.3).
  4. 4.
    And finally, the entire framework of reducing $(\Delta+1)$ coloring to extending the coloring to $O(m/\Delta)$ uncolored edges using Eulerian partition technique (in the proof of Theorem 1.1) leads to a recursion depth of $O(\log{n})$ leading to another $O(\log{n})$ overhead in the runtime.
All in all, these factors lead to our $O(m\log^{3}{n})$ time algorithm (note that the losses in Steps 1 and 2 do not compound as we run these algorithms sequentially and thus pay an additive factor of $O(\log{n})$ between the two of them, not a multiplicative one).
We now show how these log factors can be removed or replaced by $O(\log{\Delta})$ in the following. Throughout, we continue our assumption in Section 3.2 that the high probability events of the hash maps used in the algorithm happen and we ignore their randomness (e.g., if an algorithm is deterministic modulo use of hash maps, we simply refer to it as deterministic here).
Step 1: The following lemma shows the $O(\log{n})$-term of Lemma 4.1 can be easily circumvented.
Lemma A.2 (A slight modification of Lemma 4.1).
There is an algorithm that, given a graph $G$, a partial $(\Delta+1)$-edge coloring $\chi$, and $\lambda$ uncolored edges, does one of the following in $O(m+\lambda\cdot\Delta)$ time deterministically:
  1. 1.
    Extends the coloring to $\Omega(\lambda)$ uncolored edges.
  2. 2.
    Modifies $\chi$ to obtain a separable collection of $\Omega(\lambda)$ u-fans.
Proof Sketch.
The $O(\log{n})$ term of the runtime in the proof of Lemma 4.1 comes from the Harmonic series calculation in the proof of Lemma 5.13. Specifically, for any color $\alpha$, assuming we start with $\lambda_{\alpha}$ $\alpha$-primed u-edges, by 5.12, the runtime needed for “handling” the $i^{th}$ such edge is $O((m_{\alpha}+\lambda_{\alpha})/i)$ time (by handling, we mean either coloring this edge or turning it into a u-fan). In the proof of Lemma 4.1, using 5.11, this results in a total runtime of
$O(\Delta\cdot\lambda_{\alpha})+O(1)\cdot\sum_{i=1}^{\lambda_{\alpha}}\frac{m_{\alpha}+\lambda_{\alpha}}{i}=O(\Delta\cdot\lambda_{\alpha}+(m_{\alpha}+\lambda_{\alpha})\log\lambda_{\alpha})=O(\Delta\cdot\lambda_{\alpha}+(m_{\alpha}+\lambda_{\alpha})\log{n}).$
However, a simple observation is that we do not need to handle all the $\lambda_{\alpha}$ $\alpha$-primed u-edges given that our final goal is to extend the coloring or create u-fans out of only a constant fraction of the initial $\lambda$ uncolored edges. Specifically, suppose our goal is to reduce the number of $\alpha$-primed u-edges to some $\gamma\cdot\lambda_{\alpha}$ for any arbitrary constant $\gamma\in(0,1)$; then, the runtime will be
$O(\Delta\cdot\lambda_{\alpha})+O(1)\cdot\sum_{i=\gamma\cdot\lambda_{\alpha}}^{\lambda_{\alpha}}\frac{m_{\alpha}+\lambda_{\alpha}}{i}=O\left(\Delta\cdot\lambda_{\alpha}+(m_{\alpha}+\lambda_{\alpha})\cdot\log{\left(\frac{\lambda_{\alpha}}{\gamma\cdot\lambda_{\alpha}}\right)}\right)=O(\Delta\cdot\lambda_{\alpha}+m_{\alpha}),$
given that $\gamma$ is a constant. As such, the time it takes to handle any arbitrary large constant fraction of u-edges will be only $O(m)$ (by summing over choices of $\alpha\in[\Delta+1]$). The rest of the proof follows exactly as in Lemma 4.1 by simply removing the small constant fraction of u-edges not handled from the graph for the rest of the proof. ∎
Step 2: We remove the extra $O(\log{n})$-term of Lemma 4.2 in a similar way although by focusing on the expected runtime of the algorithm (plus a crucial tail inequality). In the last step, we show how to recover the final result even from this weaker guarantee.
Lemma A.3 (A slight modification of Lemma 4.2).
There is an algorithm that, given a graph $G$, a partial $(\Delta+1)$-edge coloring $\chi$ and a separable collection of $\lambda$ u-fans, extends $\chi$ to $\Omega(\lambda)$ in $T$ (randomized) time such that for some $T_{0}=O(m)$, we have,
$\mathbb{E}\left[T\right]\leq T_{0}\quad\text{and for all $\delta>0$}\quad\Pr\left[T\geq(1+\delta)\cdot 2T_{0}\right]\leq\exp\left(-\frac{\delta\cdot\lambda}{100}\right).$
Proof Sketch.
The runtime for coloring each single edge in Lemma 4.2 is $O(m/\lambda)$ by Lemma 6.7 assuming the coloring succeeds, which happens with probability at least $1/4$ by Lemma 6.5. We emphasize that this is in an average sense: the algorithm Color-U-Fans first uses Prime-U-Fans to prime $\Omega(\lambda/\Delta)$ u-fans in expected $O(m/\lambda)$ time per u-fan and thus $O((\lambda/\Delta)\cdot(m/\lambda))=O(m/\Delta)$ expected total time, and then deterministically colors them in $O(m/\Delta)$ time using Activate-U-Fans. It then repeats this process $\Delta/2$ times to color $\Omega(\lambda)$ edges, implying that $\mathbb{E}\left[T\right]=O(m)$ time.
We now prove the desired tail inequality on $T$ as well. Let us number all iterations of the while-loop in Prime-U-Fans from $1$ to $t$, across all $\Delta/2$ times Prime-U-Fans is called in Color-U-Fans: for the $i^{th}$ iteration, define an indicator random variable $X_{i}$ which is $1$ iff this iteration of the while-loop is successful in priming a u-fan (i.e., increasing the size of $\Phi$). Additionally, for every such $i$, define $S_{i}:=\sum_{j=1}^{i}X_{j}$ which denotes the number of successful iterations after the algorithm has done $i$ iterations in total. Thus, $t$ is the smallest integer where $S_{t}=\gamma\cdot\lambda$ for some integer $\gamma\in(0,1)$ which is the fraction of u-fans the algorithm colors (precisely, $\gamma=1/48\cdot 1/2\geq 1/100$).
For any $i\geq 1$, $S_{i}$ stochastically dominates the binomial distribution with parameters $i$ and $p=1/4$ (by Lemma 6.5 each iteration is successful with probability at least $1/4$). Hence, for $\delta>0$, by concentration results for the binomial distribution,
$\Pr\left[t\geq(1+\delta)\cdot 8\gamma\cdot\lambda\right]\leq\exp\left(-\frac{(2(1+\delta))^{2}}{2+2(1+\delta)}\cdot 4\gamma\cdot\lambda\right)\leq\exp\left(-\frac{\delta\cdot\lambda}{100}\right),$
using a loose upper bound in the last step. Given that the runtime of the algorithm is $t\cdot O(m/\lambda)$, the tail inequality follows. ∎
We note Lemma A.3 guarantees that as long as $\lambda=\omega(\log{n})$, the $O(m)$ runtime of the algorithm also holds with high probability. For smaller values of $\lambda$ (which happens only as a corner case in the algorithm [^15] The only case in our algorithm where this tail bound cannot be replaced with a high probability bound is when $m/\Delta=o(\log{n})$ and at the same time $\Delta\log^{2}\Delta=\omega(m\log{\Delta})$, which means $m=o(n\log{n})$ and yet $\Delta=\omega(n/\log{n})$.), we need the specific tail inequality proven in the lemma instead.
Step 3: In the next step, we provide a more fine-grained version of Lemma 4.3 that will be used in the last step of the argument to (somewhat) bypass the $O(\log{n})$ factor loss of the original lemma.
Lemma A.4 (A slight modification of Lemma 4.3).
There exists an absolute constant $b>0$ such that the following holds. There is an algorithm given a graph $G$, a partial $(\Delta+1)$-edge coloring $\chi$ of $G$ with $\lambda$ uncolored edges, and an integer $\lambda_{0}\leq\lambda$, extends $\chi$ to all but $\lambda_{0}$ uncolored edges in $T$ (randomized) time such that for some $T_{0}=O(m\cdot\log{(\lambda/\lambda_{0})}+\Delta\lambda)$, we have,
$\mathbb{E}\left[T\right]\leq T_{0}\quad\text{and for all $\delta>0$}\quad\Pr\left[T\geq(1+\delta)\cdot 2T_{0}\right]\leq\exp\left(-\frac{\delta\cdot\lambda_{0}}{200}\right).$
Proof.
The algorithm is verbatim as in Lemma 4.3 by replacing Lemma 4.1 with Lemma A.2 and Lemma 4.2 with Lemma A.3: repeatedly color a constant fraction of remaining edges as long as $\lambda_{0}$ uncolored edges remain. Since we start with $\lambda$ uncolored edges and finish with $\lambda_{0}$ many, and reduce uncolored edges by a constant factor each time, we apply Lemmas A.2 and A.3 for a total of $O(\log{(\lambda/\lambda_{0})})$ times. Moreover, the $\Delta\lambda$ terms in the runtime of Lemma A.2 form a geometric series and thus their contribution to the runtime is $O(\Delta\lambda)$ in total. This proves the expected bound on the runtime, i.e., $\mathbb{E}\left[T\right]\leq T_{0}=O(m\cdot\log{(\lambda/\lambda_{0})}+\Delta\lambda)$.
As for the tail bound, for every application of Lemma A.3 with an intermediate value $\lambda^{\prime}\in[\lambda_{0},\lambda]$, and $T^{\prime}_{0}=O(m)$ being the $T_{0}$-parameter of Lemma A.3, we have,
[…]
Thus, by union bound,
Pr[total $\Pr\left[\text{total runtime}\geq k\cdot T_{0}\right]$ $\leq\sum_{\lambda^{\prime}}\exp\left(-\frac{\delta\cdot\lambda^{\prime}}{100}\right)$
$\leq\sum_{i=0}^{\infty}\cdot\exp\left(-\frac{\delta\cdot\lambda_{0}\cdot\gamma^{i}}{100}\right)$ (as number of uncolored edges drops by some constant factor, say, $\gamma=1+\Theta(1)$ each time)
$\leq\exp\left(-\frac{\delta\cdot\lambda_{0}}{200}\right),$
by (loosely) upper bounding the sum of the geometric series using its first term. ∎
Step 4: Finally, we show how to put these lemmas together to $(\Delta+1)$ edge color the graph in $O(m\log{\Delta})$ randomized time, and conclude the proof of Theorem A.1.
Before getting to the main argument, we need the following specialized concentration inequality that we prove here (although, its formulation appears general enough to be of use elsewhere also). The proof follows a standard moment generating function argument.
Proposition A.5.
Let $\{X_{i}\}_{i=1}^{n}$ be $n$ independent non-negative random variables associated with parameters $\{\alpha_{i}\}_{i=1}^{n}$ and $\{\beta_{i}\}_{i=1}^{n}$ such that for each $i\in[n]$, $\alpha_{i},\beta_{i}\geq 1$, and for every $\delta>0$,
$\Pr\left[X_{i}\geq(1+\delta)\cdot\alpha_{i}\right]\leq\exp\left(-\delta\cdot\beta_{i}\right);$
then, for every $t\geq 0$,
$\Pr\left[\sum_{i=1}^{n}X_{i}\geq\sum_{i=1}^{n}\alpha_{i}+t\right]\leq\exp\left(-\left(\min_{i=1}^{n}\frac{\beta_{i}}{2\alpha_{i}}\right)\cdot\left(t-2\sum_{i=1}^{n}\frac{\alpha_{i}}{\beta_{i}}\right)\right).$
Proof.
Let $s:=1/2\cdot\min_{i}\beta_{i}/\alpha_{i}$ and for $i\in[n]$, $Y_{i}:=\max(X_{i}-\alpha_{i},0)$. Letting $Y:=\sum_{i=1}^{n}Y_{i}$, leads
$\Pr\left[\sum_{i=1}^{n}X_{i}\geq\sum_{i=1}^{n}\alpha_{i}+t\right]\leq\Pr\left[Y\geq t\right],$
for all $t\geq 0$. We further have,
$\Pr\left[Y\geq t\right]=\Pr\left[e^{s\cdot Y}\geq e^{s\cdot t}\right]\leq e^{-s\cdot t}\cdot\mathbb{E}\left[e^{s\cdot Y}\right]=e^{-s\cdot t}\cdot\mathbb{E}\left[\prod_{i=1}^{n}e^{s\cdot Y_{i}}\right]=e^{-s\cdot t}\cdot\prod_{i=1}^{n}\mathbb{E}\left[e^{s\cdot Y_{i}}\right]$ (2)
where the inequality is by Markov inequality and the final equality is by the independence of the variables. Now, for every $i\in[n]$,
[…]
$=\frac{1}{\beta_{i}\cdot(1-s\cdot({\alpha_{i}}/{\beta_{i}}))}$ (since the integral of $e^{-Kx}=-e^{-Kx}/K$ and $s<\beta_{i}/\alpha_{i}$ thus the value of this function in $\infty$ is zero)
$\leq\frac{1}{1-s\cdot(\alpha_{i}/\beta_{i})}$ (as $\beta_{i}\geq 1$)
$\leq\exp\left(2s\cdot\alpha_{i}/\beta_{i}\right).$ (as $s\cdot\alpha_{i}/\beta_{i}<1/2$ and $1-x\geq e^{-2x}$ for $x<1/2$)
Plugging this bounds in Equation 2, implies that
$\Pr\left[Y\geq t\right]\leq\exp\left(-s\cdot t+\sum_{i=1}^{n}2s\cdot\frac{\alpha_{i}}{\beta_{i}}\right)=\exp\left(-\left(\min_{i=1}^{n}\frac{\beta_{i}}{2\alpha_{i}}\right)\cdot\left(t-2\sum_{i=1}^{n}\frac{\alpha_{i}}{\beta_{i}}\right)\right),$
by the choice of $s$, concluding the proof. ∎
We now get to the main parts of the argument. The first part is based on a modification to the Eulerian partition approach used in the proof of Theorem 1.1.
Lemma A.6.
There is an algorithm that given a graph $G$, finds a $(\Delta+1)$ edge coloring of all but $m/\Delta$ edges in $O(m\log{\Delta})$ time with high probability.
Proof.
Consider the following recursive algorithm. Find an Eulerian tour of $G$ and partition $G$ into two edge-disjoint subgraphs $G_{1}$ and $G_{2}$ on the same vertex set such that $\Delta(G_{1}),\Delta(G_{2})\leq\lceil\Delta/2\rceil$ and $m(G_{1}),m(G_{2})\leq\lceil m/2\rceil$. For $i\in\{1,2\}$, recursively, find a $(\Delta(G_{i})+1)$-edge coloring $\chi_{i}$ of $G_{i}$ that leaves $m(G_{i})/\Delta(G_{i})=O(m/\Delta)$ edges uncolored. Combining $\chi_{1}$ and $\chi_{2}$ and uncoloring the two smallest color classes, gives a $(\Delta+1)$ coloring $\chi$ of all but $\lambda:=O(m/\Delta)$ edges. We then run Lemma A.4 to reduce the number of uncolored edges to $\lambda_{0}:=m/\Delta$ edges.
The correctness of the algorithm follows immediately. For the runtime, we have $k:=O(\Delta)$ sub-problems in total, with $2^{d}$ subproblems on $m/2^{d}$ edges and maximum degree $\Delta/2^{d}$ for $d\leq\left\lceil\log{\Delta}\right\rceil$. Let $X_{1},\ldots,X_{k}$ denote the runtime of these subproblems and thus $X:=\sum_{i=1}^{k}X_{i}$ is the total runtime. We have,
$\mathbb{E}\left[X\right]=\sum_{i}\mathbb{E}\left[X_{i}\right]=\sum_{d=1}^{\left\lceil\log{\Delta}\right\rceil}2^{d}\cdot O\left(\frac{m}{2^{d}}\log{(\lambda/\lambda_{0})}+\frac{\Delta}{2^{d}}\cdot\lambda\right)=O(m\log{\Delta})$
by Lemma A.4, since $\lambda=O(m/\Delta)$, $\lambda_{0}=m/\Delta$, and thus $\log{(\lambda/\lambda_{0})}=O(1)$ and $\Delta\cdot\lambda=O(m)$. Moreover, by Lemma A.4, for the sub-problem corresponding to $X_{i}$ at some level $d$ of the recursion, $T^{d}_{0}=O(m/2^{d})$, and every $\delta\geq 0$,
$\Pr\left[X_{i}\geq(1+\delta)\cdot 2T^{d}_{0}\right]\leq\exp\left(-\delta\cdot\frac{\lambda_{0}}{200}\right)=\exp\left(-\delta\cdot\frac{m}{200\Delta}\right).$
For this $X_{i}$, define $\alpha_{i}:=2T^{d}_{0}$ and $\beta_{i}:=m/200\Delta$. We can now apply Proposition A.5 and for $t:=1000\sum_{i=1}^{k}\alpha_{i}=1000\cdot\sum_{d=1}^{\left\lceil\log{\Delta}\right\rceil}2^{d}\cdot 2T^{d}_{0}=O(m\log{\Delta})$, have,
$\Pr\left[X\geq t\right]$ $\leq\exp\left(-\left(\min_{i=1}^{n}\frac{\beta_{i}}{2\alpha_{i}}\right)\cdot\left(t-2\sum_{i=1}^{n}\frac{\alpha_{i}}{\beta_{i}}\right)\right)$
$=\exp\left(-\frac{\Theta(1)}{\Delta}\cdot\left(t-2\cdot\frac{t\cdot 200\Delta}{1000\cdot m}\right)\right)$
$\leq\exp\left(-\frac{\Theta(1)}{\Delta}\cdot(3t/5)\right)$ (as $\Delta\leq m$ always trivially)
$=\exp\left(-\Theta(1)\cdot\frac{m\log{\Delta}}{\Delta}\right),$
where we can additionally make the constant in $\Theta(1)$ arbitrarily large without changing the asymptotic runtime of the algorithm. This probability is always at most $1/\operatorname*{poly}(n)$ because either $\Delta<\sqrt{m}$, in which case, the probability is at most $1/2^{\Theta(\sqrt{m})}$, or $\Delta>\sqrt{m}$, and thus $\log{\Delta}=\Theta(\log{n})$ and the probability is $1/\operatorname*{poly}(n)$ at the very least. This concludes the proof. ∎
In the next part, we design an algorithm that color the remaining $m/\Delta$ uncolored edges left by Lemma A.6. This can be easily done in $O(m\log{n})$ time using Lemma A.4, but doing this in $O(m\log{\Delta})$ time instead requires one more ingredient.
For this step, we use the following result from [BD23] (itself based on [Ber22, Chr23]) that allows for coloring a random uncolored edge in essentially (but not exactly) $\operatorname*{poly}(\Delta)$ time. We note that a recent work of [BD24] presents a further improvement to these bounds (in terms of the exponent of $\Delta$) but as our dependence on $\Delta$ is only logarithmic, this improvement is inconsequential and thus we stick with the simpler work of [BD23].
Proposition A.7 ([BD23]).
There is an algorithm that, given a graph $G$, a partial $(\Delta+1)$-edge coloring $\chi$, and $\lambda$ uncolored edges, extends the coloring to a randomly chosen uncolored edge in expected $\operatorname*{poly}(\Delta)\cdot\log{(n/\lambda)}$ time.
We use this result to obtain the following lemma for our purpose.
Lemma A.8.
Let $c$ be the $\operatorname*{poly}(\Delta)$-exponent in Proposition A.7, and $\lambda_{0}:=\max\{{m}/{\Delta^{2c}},m\log{n}/n\}$. There is an algorithm that, given a graph $G$, a partial $(\Delta+1)$-edge coloring $\chi$ and $\lambda=\Theta(\lambda_{0})$ uncolored edges, extends the coloring to all uncolored edges in $O(m\log\Delta)$ time with high probability.
Proof.
Suppose first that the max-term of $\lambda_{0}$ is the second one, i.e., $\lambda_{0}=m\log{n}/n$. This implies that $\Delta\geq(n/\log{n})^{1/2c}$ and since $c$ is a constant, we have $\log(\Delta)=\Theta(\log{n})$ in this case. We can use the standard Vizing’s fan and Vizing’s chain approach in Lemma 2.2 to color the remaining edges deterministically in $O(\lambda_{0}\cdot n)=O(m\log{n}/n\cdot n)=O(m\log{\Delta})$ time in this case.
Now consider the case when $\lambda_{0}=m/\Delta^{2c}$. We run the algorithm of Proposition A.7 with the following modification: when coloring a random edge, if the runtime becomes two times more than the expected time guarantee of Proposition A.7, we terminate the coloring and repeat from the beginning. Thus, each step takes $O(\Delta^{c}\cdot\log{(n/i)})$ time deterministically and succeeds in extending the coloring by one edge with probability at least half.
For $i\in[\lambda]$, define $\alpha_{i}:=4\log{(n/i)}$ and $X_{i}$ as the random variable denoting $\ell_{i}\cdot\alpha_{i}$ where $\ell_{i}$ is the number of times the algorithm attempts to color the next random edge until it succeeds when there are only $i$ edges left uncolored. We first prove that
$\sum_{i=1}^{\lambda}\alpha_{i}=o(m/\Delta^{c}),$ (3)
and then show that
$\Pr\!\left[\sum_{i=1}^{\lambda}X_{i}\geq\frac{2m}{\Delta^{c}}\right]\ll 1/\operatorname*{poly}(m),$ (4)
which concludes the proof in this case as well since the runtime of coloring the remaining edges is $O(\Delta^{c_{0}}\cdot\sum_{i=1}^{\lambda}X_{i})$ as stated earlier.
Proof of Equation 3. We have,
$\sum_{i=1}^{\lambda}\alpha_{i}$ $=\sum_{i=1}^{\lambda}4\log{\!\left(\frac{n}{i}\right)}=4\log\!\left(\prod_{i=1}^{\lambda}\frac{n}{i}\right)=4\log\!\left(\frac{n^{\lambda}}{\lambda!}\right)\leq\log\!\left(\left(\frac{e\cdot n}{\lambda}\right)^{\lambda}\right)$ (using the inequality $x!\geq(x/e)^{x}$ for all $x\geq 1$)
$=4\cdot\frac{m}{\Delta^{2c}}\cdot\log\left(\frac{e\cdot n\cdot\Delta^{2c}}{m}\right)=o(m/\Delta^{c}).$ (as $\lambda=m/\Delta^{2c}$ and $m\geq n$)
Proof of Equation 4. Note that for any $i\in[\lambda]$ and $\delta>0$,
$\Pr\left[X_{i}\geq(1+\delta)\cdot\alpha_{i}\right]=\Pr\left[\ell_{i}\geq(1+\delta)4\right]\leq 2^{-(3+4\delta)}\leq\exp\left(-\delta\right),$
where the first inequality is because the probability that $\ell_{i}$ increases by one each time is at most $1/2$ by Markov’s inequality.
While $X_{i}$’s are technically not independent (the choice of some edge being colored may change the structure of the graph for coloring next uncolored edge), they are stochastically dominated by independent random variables with above tail bounds (given the guarantee of Proposition A.7 for each uncolored edge). Thus, we can apply Proposition A.5 with parameters $\{\alpha_{i}\}_{i=1}^{\lambda}$ as above and $\beta_{i}=1$ for all $i\in[\lambda]$ to obtain that for $t:=m/\Delta^{c}$,
$\Pr\!\left[\sum_{i=1}^{\lambda}X_{i}\geq\frac{2m}{\Delta^{c}}\right]\leq\Pr\left[\sum_{i=1}^{\lambda}X_{i}\geq\sum_{i=1}^{\lambda}\alpha_{i}+t\right]$ $\leq\exp\left(-\frac{1}{4\log{n}}\cdot\frac{m}{2\Delta^{c}}\right)\ll 1/\operatorname*{poly}(m),$
where the first two inequalities use Equation (3) for $\sum_{i}\alpha_{i}$ and the third uses $\Delta^{c}\leq(n/\log{n})^{c/2c}<\sqrt{n}$ in this case (when $\lambda_{0}$ is the first-term of its maximum). This concludes the proof. ∎
Finally, we present an algorithm for taking care of the gap between the $m/\Delta$ uncolored edges left by Lemma A.6 and the $\Theta(1)\cdot\max\{m/\Delta^{2c},m\log{n}/n\}$ edges that Lemma A.8 can handle.
Lemma A.9.
Let $c=\Theta(1)$ and $\lambda_{0}:=\max\{{m}/{\Delta^{2c}},m\log{n}/n\}$ be as in Lemma A.8. There is an algorithm that, given a graph $G$, a partial $(\Delta+1)$-edge coloring $\chi$ with $\lambda=m/\Delta$ uncolored edges, extends the coloring to all but $\Theta(\lambda_{0})$ uncolored edges in $O(m\log{\Delta})$ time with high probability.
Proof.
We simply run our Lemma A.4 with the given parameters $\lambda$ and $\lambda_{0}$. By Lemma A.4, the expected runtime will be some $T_{0}=O(m\log{(\lambda/\lambda_{0})}+\Delta\lambda)=O(m\log{\Delta})$ since $\lambda_{0}\geq m/\Delta^{2c}$ for some absolute constant $c\geq 1$. Moreover, by the same lemma,
Pr[runtime of the $\Pr\left[\text{runtime of the algorithm}\geq 2000\cdot T_{0}\right]\leq\exp\left(-10\lambda_{0}\right)\leq\exp\left(-\Theta(m\log{n})/n\right)=1/\operatorname*{poly}(n),$
since $\lambda_{0}\geq m\log{n}/n$. Thus, the algorithm finishes in $O(m\log{\Delta})$ time with high probability. ∎
Putting our above lemmas together concludes the proof of Theorem A.1 as follows:
Proof of Theorem A.1.
The proof follows immediately from the lemmas in Step 4:
  • By Lemma A.6, we can color all but $m/\Delta$ edges in $O(m\log{\Delta})$ time with high probability;
  • Then, by Lemma A.9, we can extend the coloring to all but $\Theta(\lambda_{0})$ edges of the graph in $O(m\log{\Delta})$ time with high probability, where $\lambda_{0}$ is as in Lemma A.8;
  • Finally, by Lemma A.8, we can color all the remaining $\Theta(\lambda_{0})$ uncolored edges in $O(m\log{\Delta})$ time with high probability.
This concludes the proof. ∎